### v1

# Sharpness-Aware Minimization Can Hallucinate Minimizers

Abstract
Sharpness-Aware Minimization (SAM) is a widely used method that steers training toward flatter minimizers, which typically generalize better. In this work, however, we show that SAM can converge to hallucinated minimizers—points that are not minimizers of the original objective. We theoretically prove the existence of such hallucinated minimizers and establish conditions for local convergence to them. We further provide empirical evidence demonstrating that SAM can indeed converge to these points in practice. Finally, we propose a simple yet effective remedy for avoiding hallucinated minimizers.
(a) Landscape of $f$
(b) Landscape of $f^{\mathrm{SAM}}$
(c) SAM gradient field
Figure 1: Illustrative example of hallucinated minimizers. See Appendix D for details. (a) Smooth function $f$ with a minimizer set and an isolated maximizer. (b) $f^{\text{SAM}}=f(x+\rho\tfrac{\nabla f(x)}{\|\nabla f(x)\|})$; its minimizers do not correspond to minimizers or stationary points of $f$ and are therefore hallucinated. (c) Vector field of $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$; the hallucinated minimizers are attractors of the SAM iteration.
1 Introduction
It has been empirically observed in deep learning that flat minimizers tend to generalize better than sharp ones (Neyshabur et al., 2017; Jiang et al., 2020). Motivated by this observation, Sharpness-Aware Minimization (SAM) was proposed as a training method that fits the data while simultaneously regularizing against the sharpness (Foret et al., 2021; Zheng et al., 2021; Wu et al., 2020). Specifically, for a differentiable function $f\colon\mathbb{R}^{d}\to\mathbb{R}$, SAM minimizes the objective
$f^{\mathrm{SAM}}(x):=f\!\left(x+\rho\,\frac{\nabla f(x)}{\|\nabla f(x)\|}\right),$
where the perturbation radius $\rho>0$ controls the strength of the regularization. By encouraging convergence to flat minimizers, SAM has demonstrated strong empirical performance and has inspired a variety of practical variants (Foret et al., 2021).
However, although prior theoretical studies have analyzed the convergence of SAM to a minimizer under various conditions, most rely on assumptions that rarely hold in deep learning, such as convexity (Si & Yun, 2024) or a decaying perturbation radius (Khanh et al., 2024; Oikonomou & Loizou, 2025). Consequently, the validity of these convergence results is not guaranteed in practical settings.
Contributions.
In this paper, we theoretically and empirically demonstrate that SAM can, in fact, hallucinate minimizers, in the sense that its iterates may converge to points that are not minimizers of the original objective, as illustrated in Figure 1. This finding reveals a previously unrecognized failure mode of SAM in deep learning, one that is fundamentally distinct from issues caused by local minima or saddle points. To address this, we present a simple switching strategy that effectively avoids hallucinated minimizers.
1.1 Related work
SAM and its variants.
Building on the observation that flat minimizers are stable under small perturbations (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017; Chaudhari et al., 2019), SAM was proposed as a method to seek such minimizers (Foret et al., 2021). SAM has demonstrated remarkable performance across a wide range of deep learning tasks (Foret et al., 2021; Bahri et al., 2022; Zhong et al., 2022; Lee et al., 2023), motivating numerous extensions.
One line of work improves SAM by modifying its perturbation direction. Kwon et al. (2021) adapted the perturbation in a scale-invariant manner, while Kim et al. (2022) redefined it using Fisher information geometry. Li et al. (2024) removed the full-batch gradient from the perturbation direction and leveraged stochastic gradient noise for generalization. Li & Giannakis (2024) incorporated momentum into the perturbation to suppress variance and stabilize the adversary. Instead of altering the perturbation itself, Zhuang et al. (2022) adjusted the gradient update through orthogonal decomposition to reduce the surrogate gap.
Another line of research addresses the computational overhead of SAM, which stems from requiring two gradient computations per step. Some approaches reduce the number of gradient evaluations: Liu et al. (2022) computed the perturbation only periodically, Jiang et al. (2023) activated SAM only when the gradient norm is large, and Du et al. (2022b) reused gradients to avoid the second computation. Others restrict the scope of perturbations: Du et al. (2022a) applied them only to a random subset of parameters and sharpness-sensitive data, while Mueller et al. (2023) showed that limiting them to normalization layers preserves most of the benefits. Beyond reducing computation, Xie et al. (2024) improved training efficiency by parallelizing the two gradient computations.
Theoretical analyses of SAM.
Alongside its practical success, a growing body of theoretical work has analyzed SAM and examined its generalization properties from multiple perspectives. Wen et al. (2022) formalized the precise notion of sharpness minimized by SAM, clarifying its regularization effect. Möllenhoff & Khan (2023) reinterpreted SAM as a relaxation of Bayesian inference. Chen et al. (2023) showed that SAM mitigates noise fitting and improves generalization over stochastic gradient descent. Wei et al. (2023) further demonstrated that SAM alone can enhance adversarial robustness while maintaining clean accuracy.
Another line of work investigates the training dynamics and stability of SAM. Compagnoni et al. (2023) analyzed SAM through the lens of stochastic differential equations, offering a continuous-time perspective. Bartlett et al. (2023) studied quadratic objectives, showing how SAM oscillates across narrow valleys before drifting toward wider minimizers. Dai et al. (2023) demonstrated that normalization plays a key role in stabilizing SAM and ensuring robustness. Long & Bartlett (2024) extended the edge-of-stability threshold of gradient descent to SAM, showing that it depends on the gradient norm. More recently, Zhou et al. (2025) highlighted a late-phase effect, whereby SAM selects flatter minimizers when applied in the later stages of training.
Finally, several works have investigated the convergence properties of SAM in diverse settings. Andriushchenko & Flammarion (2022) proposed USAM, an unnormalized variant obtained by removing gradient normalization, and analyzed its convergence. Si & Yun (2024) provided a systematic study across convex, strongly convex, and nonconvex regimes. Khanh et al. (2024) developed a convergence analysis of SAM and its variants within the framework of inexact gradient descent. Most recently, Oikonomou & Loizou (2025) analyzed SAM and USAM within a unified framework and proved convergence under the Polyak–Łojasiewicz condition.
Hallucinated minimizers and our contribution.
Several prior studies have reported that SAM may converge to points that are not minimizers of the original loss. Kaddour et al. (2022) empirically observed that SAM can become trapped at saddle points, and Kim et al. (2023); Compagnoni et al. (2023) provided a theoretical explanation of this phenomenon using a continuous-time model of SAM. In another line of work, Bartlett et al. (2023) showed that the SAM update is equivalent to gradient descent on a surrogate function in the quadratic case, and Si & Yun (2024) proposed a virtual loss to extend this idea, although it is rigorously defined only in one dimension and lacks guarantees in higher dimensions.
In contrast, hallucinated minimizers represent a fundamentally different failure mode of SAM. They differ from saddle points in that they are not critical points of the original loss, and from surrogate-based interpretations in that they arise directly from the SAM objective. Our analysis applies to general nonconvex and high-dimensional settings and provides a rigorous characterization of hallucinated minimizers that can emerge in practical deep learning scenarios.
1.2 Preliminaries and notation
Throughout the paper, we denote $u(x):=\nabla f(x)/\|\nabla f(x)\|$ (for $x$ such that $\nabla f(x)\neq 0$). To optimize the SAM objective $f^{\mathrm{SAM}}(x)$, we require its gradient $\nabla f^{\mathrm{SAM}}(x)$. Under the standard smoothness assumption on $f$, this gradient is given by
$\nabla f^{\mathrm{SAM}}(x)=\big(I+\rho\,\nabla u(x)\big)\,\nabla f\big(x+\rho\,u(x)\big),$
where $\nabla u(x)$ denotes the Jacobian of $u(x)$. In practice, however, for computational simplicity, one does not use the exact gradient $\nabla f^{\mathrm{SAM}}$. Instead, SAM employs the “shifted” gradient $\nabla f\big(x+\rho\,u(x)\big)$. This yields the (full-batch) SAM iteration
$x_{k+1}=x_{k}-\eta_{k}\,\nabla f\big(x_{k}+\rho\,u(x_{k})\big),\qquad k=0,1,2,\dots,$
where $x_{0}\in\mathbb{R}^{d}$ is the starting point and $\eta_{0},\eta_{1},\ldots\in\mathbb{R}_{+}$ is the sequence of step sizes. The perturbation radius $\rho>0$ controls the degree of “flatness”: larger values of $\rho$ expand the neighborhood over which the loss is minimized, thereby steering the SAM iteration towards flatter minimizers. In practice, SAM is typically implemented with stochastic gradients. When clarification is needed, we refer to SAM with full-batch gradients versus SAM with stochastic gradients.
We introduce some notation. A function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is called real-analytic if its Taylor series at any point $x_{0}$ converges to $f$ on a neighborhood of $x_{0}$. For $\alpha\in\mathbb{R}$, the $\alpha$-superlevel set of $f$ is defined as $\{x:\,f(x)\geq\alpha\}$. For a set $C\subset\mathbb{R}^{d}$, we write $\partial C$ for its boundary. A set $C$ is connected if it cannot be expressed as the union of two disjoint, nonempty open sets. The distance from a point $x\in\mathbb{R}^{d}$ to a nonempty set $C\subseteq\mathbb{R}^{d}$ is $d(x,C):=\inf_{y\in C}\|x-y\|$; if $C$ is closed, the infimum is attained, and hence $d(x,C)=\min_{y\in C}\|x-y\|$. For $\delta>0$, $B_{\delta}(x)$ is the ball centered at $x$ with radius $\delta$ and the (closed) $\delta$-neighborhood of $C$ is $\mathcal{N}_{\delta}(C)=\{x:\,d(x,C)\leq\delta\}$.
2 Existence of hallucinated minimizers
In this section, we establish—under very mild assumptions—the existence of hallucinated minimizers: local minimizers of $f^{\mathrm{SAM}}$ that are not even stationary points of the original function $f$. Formally, we define hallucinated minimizers as follows:
Definition.
A point $x\in\mathbb{R}^{d}$ is a hallucinated minimizer of $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ for $\rho>0$ if $x$ is a local minimizer of $f^{\mathrm{SAM}}=f\Big(x+\rho\tfrac{\nabla f(x)}{\|\nabla f(x)\|}\Big)$ while satisfying $\nabla f(x)\neq 0$.
When the loss function $f$ is convex, hallucinated minimizers cannot arise, as shown in Theorem A.1 of Appendix A. In practical deep learning, however, the loss function $f$ is highly nonconvex, and local maximizers of $f$ can give rise to hallucinated minimizers.
2.1 Simplified existence proof with isolated maximizers
We begin with a simplified proof under the more restrictive assumption that $f$ has an isolated local maximizer, defined as a point $x^{\bullet}$ with an open neighborhood $U$ such that
$\nabla f(x)\neq 0\,\,\text{ and }\,\,f(x)<f(x^{\bullet})\quad\text{for all }x\in U\setminus\{x^{\bullet}\}.$
Theorem 2.1.
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be continuously differentiable. Assume $f$ has a global minimizer (not necessarily unique) and an isolated local maximizer. Then, a hallucinated minimizer exists for some $\rho>0$.
$C_{\varepsilon}$$\rho=\|x_{h}-x_{\star}\|$$\nabla f(x_{h})$$x_{\star}$$_{h}xx^{\bullet}$ $x_{\star}$: global minimizer $x^{\bullet}$: local maximizer $C_{\varepsilon}$: superlevel set near $x^{\bullet}$ $x_{h}$: farthest from $x_{\star}$ on $C_{\varepsilon}$ $\nabla f(x_{h})$ points toward $x_{\star}$
Figure 2: Illustration of the proof for Theorem 2.1. The point $x_{h}$ is the farthest from $x_{\star}$ among the points in $C_{\varepsilon}$. By the method of Lagrange multipliers, its gradient $\nabla f(x_{h})$ points exactly toward $x_{\star}$.
Sketch of proof.
We provide a brief sketch of the argument, with full details deferred to Appendix A.1. Figure 2 illustrates the key idea of the construction.
[…]
Next, consider $g(x)=\|x-x_{\star}\|^{2}$ and define
$x_{h}\in\operatorname*{arg\,max}_{x\in C_{\varepsilon}}g(x)=\operatorname*{arg\,max}_{x\in\partial C_{\varepsilon}}g(x).$
In words, if $x_{h}$ maximizes $g$ over $C_{\varepsilon}$, then it must lie on the boundary $\partial C_{\varepsilon}$. Since $\partial C_{\varepsilon}$ coincides with $\{x:f(x)=f(x^{\bullet})-\varepsilon\}$ near $x_{h}$, we apply the method of Lagrange multipliers to $x_{h}\in\operatorname*{arg\,max}_{x\in\partial C_{\varepsilon}}g(x)$ to obtain
$2(x_{\star}-x_{h})=\nabla g(x_{h})=\lambda\nabla f(x_{h}).$
The fact that $\lambda>0$ follows from the observation that $\nabla f(x_{h})$ points “toward” both $x_{\star}$ and $x^{\bullet}$, as illustrated in Figure 2. Finally, setting $\rho=\lambda\|\nabla f(x_{h})\|/2$ yields
$x_{h}+\rho\tfrac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{\star},$
and hence $f^{\mathrm{SAM}}(x_{h})=f(x_{\star})$. Thus, $x_{h}$ is a (global) minimizer of $f^{\mathrm{SAM}}$. ∎
Importantly, the existence of a hallucinated minimizer also holds when $x_{\star}$ is a local minimizer, provided that $f$ has a locally Lipschitz gradient. See Appendix A.2 for details.
The proof of Theorem 2.1 reveals the conditions under which a hallucinated minimizer is likely to arise. Figure 2 illustrates the core idea: given a minimizer $x_{\star}$, $x_{h}$ satisfies $x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$. This means that at $x_{h}$, the gradient points directly toward $x_{\star}$. However, near a minimizer, the gradient typically points outward, making it difficult to identify such $x_{h}$ in its immediate neighborhood. To resolve this, Theorem 2.1 requires the presence of a nearby maximizer, which allows the gradient to align in the desired direction and thus necessitates nonconvexity of the objective.
The proof also shows that the perturbation radius $\rho$ must exceed the distance between minimizer and the maximizer, since $x_{h}$ is the farthest point from $x_{\star}$ on the superlevel set $C_{\varepsilon}$. Because $x_{h}\in C_{\varepsilon}$, a hallucinated minimizer is located near the maximizer. This implies that hallucinated minimizers are typically associated with high loss values.
Taken together, these observations suggest that hallucinated minimizers generally arise in nonconvex objectives containing both local maximizers and minimizers, and that they tend to occur near a local maximizer within a high-loss region.
2.2 Existence of hallucinated minimizers for neural networks
Theorem 2.1 assumes that the local maximizer is isolated. We now relax this assumption, since in neural networks, maximizers (like minimizers) often occur as sets.
To this end, we leverage the real-analyticity of neural networks, an assumption that holds when training on a finite dataset (empirical loss) with real-analytic activation functions. We note that all commonly used activation functions are real-analytic, except for ReLU. Technically, real-analyticity affords us the Łojasiewicz inequality, which we use to rule out certain pathological cases.
Notably, our result does not rely on restrictive but common structural assumptions on the loss function, such as global smoothness or a quadratic form.
Definition.
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function. A nonempty connected set $X^{\bullet}\subseteq\mathbb{R}^{d}$ is a local maximizer set of $f$ if there exists $\delta>0$ such that $X^{\bullet}$ is the maximizer set over its $\delta$-neighborhood. In other words,
$X^{\bullet}=\operatorname*{arg\,max}_{y\in\mathcal{N}_{\delta}(X^{\bullet})}f(y).$
Furthermore, we denote by $f(X^{\bullet})$ the common function value of $f$ on $X^{\bullet}$.
Figure 3: Illustration of a local maximizer set. The dotted unit circle is a local maximizer set of a real-analytic function and is not a singleton.
Theorem 2.2.
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be real-analytic. Assume $f$ has a global minimizer (not necessarily unique) and a bounded local maximizer set. Then, a hallucinated minimizer exists for some $\rho>0$.
The proof of Theorem 2.2, fully presented in Appendix A.4, is analogous to that of Theorem 2.1, except that the isolatedness assumption is replaced by the real-analyticity assumption. A key technical point is to rule out the possibility that critical points accumulate densely around $X^{\bullet}$. This is achieved using the Łojasiewicz inequality:
Lemma 2.3 (Łojasiewicz (1963)).
If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is real-analytic, then for every $p\in\mathbb{R}^{d}$, there exist an open neighborhood $U$ of $p$, a constant $C>0$, and an exponent $q\in(0,1)$ such that
$|f(p)-f(x)|^{q}\leq C\|\nabla f(x)\|\quad\textit{ for all }\,x\in U.$
Most modern neural networks are real-analytic.
The real-analyticity assumption on $f$ is practical in the context of deep learning. In particular, consider the empirical loss function
$f(\theta)=\sum_{i=1}^{N}\ell(h_{\theta}(X_{i}),Y_{i}),$
where $\theta\in\mathbb{R}^{d}$ denotes the neural network parameters. If the dataset is finite ($N<\infty$), $\ell(\cdot,\cdot)$ is real-analytic in its first argument (as is the case for most commonly used losses, such as cross-entropy), and $h_{\theta}$ is a neural network built from real-analytic activation functions (e.g., tanh, ELU, GELU, SiLU, swish), then $f(\theta)$ is real-analytic. More concretely, $h_{\theta}$ may be a finite composition of linear layers, convolution, attention, residual connections, layer normalization, batch normalization, real-analytic activation functions, the softmax function, average pooling, and dropout. However, $h_{\theta}$ cannot incorporate ReLU, leaky ReLU, or max-pooling, since these operations are non-smooth and therefore non-analytic.
Further discussion and details are provided in Appendix A.3.
3 Geometric and dynamical properties of hallucinated minimizers
In this section, we establish a finer geometric property of hallucinated minimizers as well as a dynamical property of the SAM iterates. Recall that we use the notation $u(x)=\nabla f(x)/\|\nabla f(x)\|$.
3.1 The set of hallucinated minimizers can have manifold structures
The following theorem establishes that when the set of true minimizers has an $m$-dimensional manifold structure, the set of hallucinated minimizers inherits the same geometric structure. In particular, this result explains why, in Figure 1, the set of SAM minimizers appears as a curve.
Theorem 3.1.
Suppose $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfies the assumptions of Theorem 2.2. Assume $\mathcal{M}\subseteq\mathrm{argmin}f$, where $\mathcal{M}\subseteq\mathbb{R}^{d}$ is a nonempty smooth $m$-dimensional manifold. Let $x_{h}$ be a hallucinated minimizer with a corresponding $\rho>0$ as constructed in the proof of Theorem 2.2. If $I+\rho\nabla u(x_{h})\in\mathbb{R}^{d\times d}$ is nonsingular, then the set of hallucinated minimizers contains a smooth manifold of dimension $m$.
We defer the proof to Appendix B.1.
3.2 Hallucinated minimizers are attractors
We have established the existence of hallucinated minimizers, but does SAM actually converge to them? Recall that the SAM iteration is given by
$x_{k+1}=x_{k}-\eta_{k}\nabla f\!\left(x_{k}+\rho\,u(x_{k})\right),\qquad k=0,1,2,\dots,$
where $\eta_{k}>0$ denotes the step size.
The answer is yes—hallucinated minimizers can indeed be attractors of the SAM dynamics. Thus, the concern about hallucinated minimizers in neural network training is not merely hypothetical. In Section 4, we provide empirical evidence that SAM can converge to hallucinated minimizers. In this subsection, we theoretically establish local convergence to hallucinated minimizers.
Theorem 3.2.
Suppose $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is real-analytic, and let $H\subset\mathbb{R}^{d}$ be a bounded, connected set of hallucinated minimizers of $f$ for a fixed perturbation radius $\rho>0$. Assume there exists $\delta>0$ such that the $\delta$-neighborhood of $H$ contains no minimizers of $f^{\mathrm{SAM}}$ other than those already in $H$. Assume further that every $x_{h}\in H$ satisfies
$1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0,\,\,\textup{ where }\,\mathrm{Sym}(\nabla u(x_{h}))=\frac{1}{2}(\nabla u(x_{h})+\nabla u(x_{h})^{\top}).$
If the initialization $x_{0}$ is chosen sufficiently close to $H$, then there exists a sufficiently small fixed step size $\eta_{k}=\eta>0$ such that the SAM iterates converge to $H$, in the sense that $d(x_{k},H)\to 0$.
We defer the proof to Appendix B.2.
4 Empirical analyses in deep learning
In this section, we empirically validate our theory by analyzing hallucinated minimizers in deep learning. We show that SAM trajectories can, in practice, converge to hallucinated minimizers. We further demonstrate that a simple switching strategy can effectively prevent this convergence, providing a practical safeguard for SAM against this failure mode.
(a) Original loss $f$ around $x_{h}$
(b) SAM loss $f^{\mathrm{SAM}}$ around $x_{h}$
(c) SAM iterates converge back to the hallucinated minimizer set.
Figure 4: Visualizations of $f$ and $f^{\mathrm{SAM}}$ around the hallucinated minimizer $x_{h}$. Plots (a) and (b) are taken on a 2-dimensional plane defined by $x_{h}$ and two random directions. These show that $x_{h}$ is indeed a minimizer of $f^{\mathrm{SAM}}$ but not a stationary point of $f$. Plot (c) depicts $f^{\mathrm{SAM}}$ on the 2-dimensional plane containing $x_{h}$, $x_{0}$, and $x_{N}$, where $x_{0}$ is a small perturbation of $x_{h}$ and $x_{N}$ is obtained after $N=1000$ SAM steps from $x_{0}$. The pink horizontal line segment indicates the set of hallucinated minimizers, showing that the SAM trajectory converges back to this set.
4.1 SAM can converge to hallucinated minimizers
We first present an example in which SAM converges to a non-minimizer and confirm that the point is a hallucinated minimizer. We then examine how frequently such convergence occurs across diverse experimental settings.
Do we really converge to hallucinated minimizers?
We begin with a simple neural network setting and provide an example where SAM converges to a hallucinated minimizer. Specifically, we train a two-layer neural network with Tanh activations on MNIST (LeCun et al., 1998) using full-batch updates, yielding a smooth objective. The experimental setup is described in Appendix C. With a perturbation radius $\rho=1.8$, we train for over 20 million steps and observe that the trajectory converges to a single point $x_{h}$. At this point, the SAM gradient nearly vanishes ($\|\nabla f(x_{h}+\rho\,u(x_{h}))\|=4.8\times 10^{-9}$), while the original gradient remains relatively large ($\|\nabla f(x_{h})\|=0.0627$). Thus, $x_{h}$ is a stationary point of $f^{\mathrm{SAM}}$ but not of $f$, showing that SAM can converge to a hallucinated stationary point.
To verify that $x_{h}$ is indeed a hallucinated minimizer, we visualize the loss landscape around it, as shown in Figure 4. Following the method of Li et al. (2018), we define the plane
$x(\alpha,\beta)=x_{h}+\alpha u+\beta v,$
where $u,v\in\mathbb{R}^{d}$ are orthogonal vectors of equal norm. In Figure 4(a), $x_{h}$ is clearly not a minimizer of $f$, whereas in Figure 4(b), on the same plane, it appears as a minimizer of $f^{\mathrm{SAM}}$. This demonstrates that SAM can converge to a hallucinated minimizer in neural networks.
To investigate the geometry around $x_{h}$ in more detail, we add a small random perturbation to $x_{h}$, yielding a nearby point $x_{0}$ with $\|x_{0}-x_{h}\|=0.1$. Starting from $x_{0}$, we run $N=1000$ additional SAM steps to reach $x_{N}$. Figure 4(c) shows the SAM objective on the plane spanned by $x_{h}$, $x_{0}$, $x_{N}$. The SAM trajectory (blue) is projected onto this plane, while the SAM minimizers (pink) are computed within a tolerance of $10^{-9}$. We observe that hallucinated minimizers form a connected set, consistent with prior work on connected minimizers in neural networks (Garipov et al., 2018) and with Theorem 3.1, which guarantees that this structure is preserved.
$\rho=1.0$
$\rho=1.3$
$\rho=1.6$
$\rho=1.9$
(a) SAM-only
(b) GD $\rightarrow$ SAM switching
Figure 5: Training loss and test accuracy for SAM on MNIST with a 2-layer network using full-batch gradients. The top row shows the results from SAM-only training, while the bottom row shows the results from the switching strategy as described in Section 4.2. Final average test accuracies are 96.77%, 90.85%, 67.88%, and 13.36% for SAM-only training, compared to 97.77%, 97.73%, 97.69%, and 97.62% for the switching strategy. These results demonstrate that the switching strategy improves test accuracy and stabilizes outcomes across perturbation radii.
How common are hallucinated minimizers?
Next, we investigate convergence to hallucinated minimizers across multiple experiments, showing that this phenomenon can occur frequently in deep learning. Figure 5 presents SAM training outcomes in the same full-batch setting as before, with 80 distinct seeds and 100,000 iterations for perturbation radii $\rho=1.0,1.3,1.6,1.9$. The top row shows SAM-only results, while the bottom row corresponds to the switching strategy discussed later. In this subsection, however, we focus on the SAM-only setting.
The experiments reveal that convergence to hallucinated minimizers depends on the perturbation radius $\rho$. At $\rho=1.0$, training consistently converges to zero loss, indicating convergence to a true minimizer of the original loss. At $\rho=1.3$ and $1.6$, some trajectories stabilize at nonzero-loss points. At $\rho=1.9$, most trajectories converge to such nonzero-loss points, indicating that SAM predominantly reaches hallucinated minimizers.
These findings provide strong evidence that hallucinated minimizers are not rare anomalies but occur consistently in deep learning when the perturbation radius $\rho$ is large. This observation aligns with Theorem 2.1, whose proof requires the perturbation radius $\rho$ to exceed the distance between a minimizer and a maximizer. At the same time, even for the same $\rho$, trajectories may converge either to a hallucinated minimizer or to a true minimizer, consistent with Theorem 3.2, which depends on the initialization’s proximity to a hallucinated minimizer.
In the stochastic case, we train ResNet-18 (He et al., 2016) on CIFAR-100 (Krizhevsky, 2009) with mini-batch updates, following the FSAM implementation of Li et al. (2024). We observe a similar trend: larger perturbation radii lead to unstable training. Further details of the experimental settings and results are provided in Appendix C.
Figure 6: Final test accuracy for SAM-only and the switching strategy on CIFAR-100 with ResNet-18 using stochastic gradients. Each curve shows the mean (bold) and standard deviation (shaded area) over 5 seeds, evaluated at perturbation radii $\rho=0.1,0.2,\ldots,0.8$. Both methods achieve peak accuracy at $\rho=0.3$, with 77.05% for SAM-only and 77.49% for the switching strategy.
4.2 Switching strategy for avoiding hallucinated minimizers
One obvious approach to avoiding hallucinated minimizers is to use a small perturbation radius $\rho>0$. Indeed, hallucinated minimizers do not arise when $\rho=0$ (i.e., when SAM is not applied). However, the perturbation itself is the key mechanism that regularizes against sharpness, making it desirable to use a moderately large value of $\rho$.
In this subsection, we introduce a simple yet effective heuristic for avoiding hallucinated minimizers, which we call switching. The idea is to use plain gradient descent for the first 10% of training iterations and then switch to SAM.
As shown in Figure 5(b), the switching strategy consistently drives the training loss to zero across all tested perturbation radii, including large values of $\rho$ for which standard SAM fails to converge to the true minima. Figure 6 further demonstrates the improved test accuracy achieved under the switching strategy. Not only does switching yield higher test accuracy, but it also reduces sensitivity to the choice of $\rho$. This is a notable advantage, since the perturbation radius $\rho$ is a difficult hyperparameter to tune in SAM.
The effectiveness of switching can be understood within our theoretical framework. The proof of Theorem 2.1 suggests that hallucinated minimizers tend to arise “near” local maximizers, and hence are more likely to occur in high-loss regions. This further implies that during the early stages of training, when loss values are large, SAM is particularly at risk of entering a basin of attraction of hallucinated minimizers. By first applying gradient descent to escape from high-loss regions before switching to SAM, this risk is mitigated. Consequently, our simple remedy ensures that SAM remains stable even for large perturbation radii.
A related mechanism was studied by Zhou et al. (2025), who applied SAM only in the later stages of training to improve generalization. In contrast, our results show that the switching strategy specifically prevents convergence to hallucinated minimizers, providing a complementary explanation for its effectiveness.
5 Conclusion
In this work, we identify a previously unrecognized failure mode of SAM: its tendency to converge to hallucinated minimizers. Our theoretical analysis establishes the existence of such minimizers under practical assumptions, and our empirical results validate the theory by demonstrating that SAM can indeed converge to them. To address this, we present a simple switching strategy that effectively avoids hallucinated minimizers.
Although our theoretical and empirical findings are consistent, gaps remain between the theoretical characterizations and broader empirical findings. These gaps open several interesting avenues for follow-up work.
One direction is to extend our theoretical analysis to the setting where SAM employs stochastic gradients rather than full-batch gradients. While our experiments suggest that hallucinated minimizers also arise in the stochastic case, a more rigorous theoretical understanding is desirable. Another direction is to analyze, from a theoretical standpoint, how likely it is for SAM to converge to hallucinated minimizers. Our experiments show that convergence to such minimizers is common, whereas our current theory only guarantees convergence within a local neighborhood of these points. Yet another direction is to extend the analysis to other variants of SAM. Our current results rely on the normalization of the ascent direction, which renders the magnitude of $\nabla f(x)$ irrelevant in constructing hallucinated minimizers. For SAM variants that incorporate gradient magnitude in the ascent step, a modified analysis would be necessary.
Reproducibility statement
We have taken extensive measures to ensure reproducibility. Complete proofs of all theorems, together with detailed assumptions, are provided in Appendix A and Appendix B. Experimental setups, including datasets and hyperparameters, are described in Appendices C and D. The implementation of our main experiments is available through a repository at https://anonymous.4open.science/r/SAM-can-hallucinate-minimizers-4B82.
Appendix A Omitted details for Section 2
We begin by showing that the hallucinated minimizers cannot exist when the loss function is convex. For ease of exposition, throughout the appendices we use the following notation: for $x\in\mathbb{R}^{d}$, let
$x^{+}:=x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}.$
Proposition A.1.
If the function $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex, then there is no point $x\in\mathbb{R}^{d}$ such that $\nabla f\left(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}\right)=0$.
Proof.
Suppose $\nabla f(x)\neq 0$ and $\nabla f(x^{+})=0$. Since $f$ is convex, $x^{+}$ must be a global minimizer. However, convexity also implies
$f(x^{+})\geq f(x)+\left\langle\nabla f(x),\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}\right\rangle=f(x)+\rho\|\nabla f(x)\|>f(x),$
which contradicts the optimality of $x^{+}$. ∎
A.1 Full proof of Theorem 2.1
We now provide the full proof of the existence theorem.
See 2.1
Proof.
Let $x^{\bullet}$ be a local maximizer with $f(x^{\bullet})=M$, and let $C\subseteq U$ be a compact ball centered at $x^{\bullet}$ such that $f(x)<f(x^{\bullet})$ for all $x\in C$. The strict inequality follows from $x^{\bullet}$ being an isolated critical point. Define $m:=\max_{\partial C}f(x)<M$ and consider the preimage $f^{-1}([M-\varepsilon,M])$ where $0<\varepsilon<M-m$. Let $C_{\varepsilon}$ denote the connected component of this preimage containing $x^{\bullet}\in f^{-1}([M-\varepsilon,M])$.
By Lemma A.2, $C_{\varepsilon}\subseteq\mathrm{int}\,\,C$, and hence $C_{\varepsilon}$ is compact. Moreover, by Lemma A.3, the function value on the boundary satisfies $f(x)=M-\varepsilon$ for all $x\in\partial C_{\varepsilon}$.
Consider the squared distance function $g(x)=\|x-x_{\star}\|^{2}$, and let
$x_{h}\in\mathrm{argmax}_{C_{\varepsilon}}g(x),\quad x_{h}\neq x_{\star},\quad\textup{ and}\quad\rho:=\|x_{h}-x_{\star}\|.$
Then, $x_{h}$ must be on the boundary of $C_{\varepsilon}$, and thus $f(x_{h})=M-\varepsilon$. Furthermore, since $x^{\bullet}\neq x_{h}$ is the only critical point of $f$ in $C_{\varepsilon}$, the gradient at $x_{h}$ does not vanish. Consequently, there exists an open neighborhood $V$ of $x_{h}$ such that $\Sigma:=\{x\in V:f(x)=M-\varepsilon\}$ is an embedded $C^{1}$ hypersurface near $x_{h}$. By shrinking $V$ if necessary, we may assume $V\cap\partial C_{\varepsilon}=\Sigma$. Thus, maximizing $g(x)$ over $C_{\varepsilon}$ is locally equivalent to maximizing $g(x)$ over the hypersurface $\Sigma$. By the method of Lagrange multipliers, we obtain
$\nabla g(x_{h})=\lambda\nabla f(x_{h}).$
That is, there exists $\lambda>0$ such that
$2(x_{\star}-x_{h})=\lambda\nabla f(x_{h})$
by Lemma A.4. Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
$x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|},$
which implies that $x_{h}$ is a hallucinated minimizer. ∎
We now prove the three lemmas used in the proof of Theorem 2.1.
Lemma A.2.
The set $C_{\varepsilon}$ from Theorem 2.1 is contained in $\mathrm{int}\,C$. Hence, it is compact.
Proof.
Suppose $x\in C_{\varepsilon}$. If $x\in\partial C$, then $f(x)\geq M-\varepsilon>m$, contradicting the definition of $m$ as the maximum value of $f$ on $\partial C$.
If instead $x\in\mathrm{ext}\,C=\mathbb{R}^{d}\setminus C$, then $\mathrm{int}\,C$ and $\mathbb{R}^{d}\setminus C$ are two nonempty disjoint open sets that separate $C_{\varepsilon}$, contradicting the fact that $C_{\varepsilon}$ is connected. Therefore, $x$ must lie in $\mathrm{int}\,C$, and $C_{\varepsilon}\subseteq\mathrm{int}\,C$. ∎
The following lemma shows that every point on $\partial C_{\varepsilon}$ in Theorem 2.1 lies on the same level set.
Lemma A.3.
Any point $x\in\partial C_{\varepsilon}$ from Theorem 2.1 satisfies $f(x)=M-\varepsilon$.
Proof.
Take $x\in\partial C_{\varepsilon}$. If $f(x)>M-\varepsilon$, continuity of $f$ implies that there exists $r>0$ such that $f(y)>M-\varepsilon$ for all $y$ in an open ball centered at $x$ with radius $r$. This contradicts the fact that $x$ is a boundary point of $C_{\varepsilon}$.
If $f(x)<M-\varepsilon$, this directly contradicts the fact that $x\in f^{-1}([M-\varepsilon,M])$. Therefore, $f(x)=M-\varepsilon$. ∎
The next lemma establishes that $\lambda>0$ in Theorem 2.1.
Lemma A.4.
In the proof of Theorem 2.1, the vectors $x_{\star}-x_{h}$ and $\nabla f(x_{h})$ point in the same direction. Equivalently, $\lambda>0$.
Proof.
Let $V$ be an open neighborhood of $x_{h}$. By possibly shrinking $V$, we may assume the local superlevel set $\{x\in V:f(x)\geq M-\varepsilon\}$ is contained in $C_{\varepsilon}$. Then, $x_{h}$ is a maximizer of $g$ over the feasible region $\{x\in V:f(x)\geq M-\varepsilon\}$.
Consider any feasible direction $d\in\mathbb{R}^{d}$ with $\langle\nabla f(x_{h}),d\rangle\geq 0$. Let $\gamma\colon(-\varepsilon,\varepsilon)\rightarrow\mathbb{R}^{d}$ be a smooth curve such that $\gamma(0)=x_{h}$, $\gamma^{\prime}(0)=d$, and $\gamma(t)\in V$. Then, $f(\gamma(t))\geq M-\varepsilon$ for sufficiently small $t>0$. Moreover,
$\frac{d}{dt}f(\gamma(t))\Bigg|_{t=0}=\langle\nabla f(x_{h}),d\rangle\geq 0,$
so $\gamma(t)$ remains in the feasible set for small $t>0$. Since $x_{h}$ maximizes $g$, it follows that
$\lim_{t\downarrow 0}\frac{d}{dt}g(\gamma(t))=\lim_{t\downarrow 0}\frac{d}{dt}\|\gamma(t)-x_{\star}\|^{2}=\lim_{t\downarrow 0}2\langle\gamma(t)-x_{\star},\gamma^{\prime}(t)\rangle=2\langle x_{h}-x_{\star},d\rangle\leq 0.$
Taking the feasible direction $d=\nabla f(x_{h})$ yields
$2\langle x_{h}-x_{\star},\nabla f(x_{h})\rangle=-2\lambda\|\nabla f(x_{h})\|^{2}\leq 0.$
Since $\lambda\neq 0$, this inequality implies $\lambda>0$. ∎
A.2 Extending Theorem 2.1 to local minimizers
We now extend Theorem 2.1 by relaxing the assumption that $x_{\star}$ is a global minimizer. In fact, a hallucinated minimizer can still exist when $x_{\star}$ is only a local minimizer. To show this, we first establish the following lemma.
Lemma A.5.
Suppose $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|$ for some $L>0$, and that $\nabla f(x)\neq 0$ and $\nabla f(y)\neq 0$. Then, we have
$\|y+\rho\,u(y)-x-\,\rho\,u(x)\|\leq\Big(1+\frac{2\rho L}{\|\nabla f(x)\|}\Big)\|y-x\|.$
Proof.
It follows from the triangle inequality that
$\begin{split}&\|y+\rho\,u(y)-x-\rho\,u(x)\|\\ &\leq\|y-x\|+\rho\Big\|\frac{\nabla f(y)}{\|\nabla f(y)\|}-\frac{\nabla f(x)}{\|\nabla f(x)\|}\Big\|\\ &\leq\|y-x\|+\rho\Big\|\frac{\nabla f(y)}{\|\nabla f(y)\|}-\frac{\nabla f(y)}{\|\nabla f(x)\|}\Big\|+\rho\Big\|\frac{\nabla f(y)}{\|\nabla f(x)\|}-\frac{\nabla f(x)}{\|\nabla f(x)\|}\Big\|\\ &\leq\|y-x\|+\rho\frac{\|\nabla f(x)-\nabla f(y)\|}{\|\nabla f(x)\|}+\rho\frac{\|\nabla f(y)-\nabla f(x)\|}{\|\nabla f(x)\|}\\ &\leq\|y-x\|\Big(1+\frac{2\rho L}{\|\nabla f(x)\|}\Big).\end{split}$
Finally, we obtain the following corollary.
Corollary A.6.
Suppose $f$ has a locally Lipschitz gradient. Then, Theorem 2.1 remains valid even when $x_{\star}$ is a local minimizer of $f$.
Proof.
Recall the last equation of the proof of Theorem 2.1:
$x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}.$
Consider an open ball centered at $x_{h}$ with radius $r$ chosen sufficiently small so that $\nabla f$ does not vanish on the ball. Let $L>0$ be such that $\|\nabla f(x_{h})-\nabla f(y)\|\leq L\|x_{h}-y\|$ for any $y\in B_{r}(x_{h})$. Since $x_{\star}$ is a local minimizer, there exists $\delta>0$ such that
$f(y)\geq f(x_{\star})\quad\forall y\mbox{ with }\|y-x_{\star}\|\leq\delta.$
Now consider an open ball centered at $x_{h}$ with radius
$r^{\star}:=\min\Big\{\frac{\delta}{1+\frac{2\rho L}{\|\nabla f(x_{h})\|}},r\Big\}.$
Then, for any $y\in B_{r^{\star}}(x_{h})$, we have
$\|x_{\star}-y-\rho\,u(y)\|=\|x_{h}+\rho\,u(x_{h})-y-\rho\,u(y)\|\leq\|x_{h}-y\|\left(1+\frac{2\rho L}{\|\nabla f(x_{h})\|}\right)\leq\delta,$
where the first inequality follows from Lemma A.5. Hence,
$f(y+\rho\,u(y))\geq f(x_{\star}).$
This implies $f^{\mathrm{SAM}}(y)\geq f^{\mathrm{SAM}}(x_{h})$, so $x_{h}$ is a local minimizer of $f^{\mathrm{SAM}}$, and therefore a hallucinated minimizer. ∎
A.3 Most modern neural networks are real-analytic: Further discussion
We now formalize the claim that the neural network $\theta\mapsto h_{\theta}(X)$ is real-analytic in its parameters $\theta$. By standard arguments, the following two lemmas establish that the final loss function
$f(\theta)=\sum_{i=1}^{N}\ell(h_{\theta}(X_{i}),Y_{i})$
is real-analytic under the assumptions on $h_{\theta}$ stated in Section 2.
Lemma A.7.
If $h_{\theta}\colon\mathbb{R}^{d}\to\mathbb{R}^{h}$ is a neural network with real-analytic activation functions, then $h_{\theta}$ is real-analytic as a function of $\theta$.
Proof.
Fix input data $x$ and let $z$ denote the hidden states, which depend on $x$ and $\theta$. Then, $z$, hence the entire network $h_{\theta}$, is a finite composition of the following real-analytic mappings:
  • (Layer normalization) $z\mapsto\gamma\odot\frac{z-\mu_{L}(z)}{\sqrt{\sigma_{L}(z)+\varepsilon}}+\beta$, where $\mu_{L}{(z)}$ and $\sigma_{L}(z)$ are the per-sample mean and variance;
  • (Batch normalization) $z\mapsto\gamma\odot\frac{z-\mu_{B}{(z)}}{\sqrt{\sigma_{B}(z)+\varepsilon}}+\beta$, where $\mu_{B}{(z)}$ and $\sigma_{B}(z)$ are the per-channel mean and variance;
  • (Activation) $z\mapsto\sigma(z)$, where $\sigma$ is a real-analytic activation function;
  • (Softmax) $z\mapsto\mu(z)$, where $\mu$ is the softmax function;
  • (Average pooling) $z\mapsto Az$, where $A$ is a linear averaging operator;
  • (Residual connection) $z\mapsto z+F_{\theta}(z)$, where $F$ is real-analytic in $\theta$;
  • (Convolution layer) $(z,W_{i},b_{i})\mapsto W_{i}*z+b_{i}$, where * denotes the discrete convolution operator;
  • (Dropout) $z\mapsto m\odot z$, where $m$ is a masking operator;
  • (Linear layer) $(z,A_{i},b_{i})\mapsto A_{i}z+b_{i}$; and
  • (Attention layer) $(z,W_{Q},W_{K},W_{V})\mapsto\mu\big(\frac{{Q}{K}^{\top}}{\sqrt{d_{\mathrm{attn}}}}\big){V}$, where $Q=zW_{Q}$, $K=zW_{K}$, $V=zW_{V}$, and $d_{\mathrm{attn}}$ is the size of the attention matrix $Q,K$.
Since the composition of real-analytic functions is real-analytic, it follows that $h_{\theta}$ is real-analytic. ∎
Finally, the next lemma establishes that $f$ is real-analytic. This follows directly from the fact that the composition of real-analytic functions is real-analytic; hence the proof is omitted.
Lemma A.8.
Let $h_{\theta}$ be a (finite) neural network constructed with linear layers, attention, convolution, layer normalization, and real-analytic activation functions (all commonly used activation functions except ReLU are real-analytic). Then, for a real-analytic loss function $\ell$,
$f(\theta):=\frac{1}{N}\sum^{N}_{i=1}\ell(h_{\theta}(X_{i}),Y_{i})$
is real-analytic as a function of $\theta$.
A.4 Full proof of Theorem 2.2
The following lemma shows that, under the real-analyticity assumption, critical points cannot accumulate around the local maximizer set $X$. The argument relies on the Łojasiewicz inequality.
Lemma A.9.
Suppose $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is real-analytic and $X$ is a bounded local maximizer set of $f$ with some $\delta>0$. Then, there exists $\varepsilon>0$ with the following property: if $x$ is a critical point in the $\delta$-neighborhood of $X$ with $f(x)\geq f(X)-\varepsilon$, then $x\in X$.
Proof.
Define the closed $\delta$-neighborhood of $X$ by $\mathcal{N}_{\delta}(X):=\{y:d(y,X)\leq\delta\}$. Since $X$ is a bounded connected set, $\mathcal{N}_{\delta}(X)$ is compact and connected. Let $S$ denote the set of critical points in $\mathcal{N}_{\delta}(X)$ that are not in $X$:
$S=\{s\in\mathcal{N}_{\delta}(X):\nabla f(s)=0\}\setminus X.$
If $S=\emptyset$, then the theorem holds for any $\varepsilon>0$, and we are done. Assume instead that $S\neq\emptyset$. For each point $x\in X$, Lemma 2.3 guarantees the existence of an open neighborhood $U_{x}$, a constant $C_{x}>0$, and an exponent $q_{x}\in(0,1)$ such that
$|f(x)-f(y)|^{q_{x}}=|f(X)-f(y)|^{q_{x}}\leq C_{x}\|\nabla f(y)\|,\,\quad\,y\in U_{x}.$
If $y\in S$, then
$|f(X)-f(y)|^{q_{x}}\leq C_{x}\|\nabla f(y)\|=0.$
This implies $f(y)=f(X)$, which is a contradiction. Hence, $S\subseteq C:=\mathcal{N}_{\delta}(X)\setminus\bigcup_{x\in X}U_{x}$. (Note that $C\neq\emptyset$ since $S\neq\emptyset$). Since $C$ is compact, define
$0<\varepsilon^{\star}:=f(X)-\max_{y\in C}f(y).$
Then, any $\varepsilon\in(0,\varepsilon^{\star})$ satisfies the theorem. ∎
Finally, we use Lemma A.9 to complete the proof of Theorem 2.2.
See 2.2
Proof.
Define the closed $\delta$-neighborhood of a local maximizer set $X$ as $\mathcal{N}_{\delta}(X):=\{y:d(y,X)\leq\delta\}$. Since $X$ is a bounded connected set, $\mathcal{N}_{\delta}(X)$ is compact and connected. By Lemma A.9, there exists $\varepsilon_{1}>0$ such that any critical point in $\mathcal{N}_{\delta}(X)$ with function value at least $f(X)-\varepsilon_{1}$ must lie in $X$. Next, choose $\varepsilon_{2}>0$ such that $0<\varepsilon_{2}<f(X)-\max_{\partial\mathcal{N}_{\delta}(X)}f(x)$.
Let $\varepsilon:=\min\{\varepsilon_{1},\varepsilon_{2}\}$, and consider the preimage $f^{-1}([f(X)-\varepsilon,f(X)])$. Define $C_{\varepsilon}$ as the connected component of this preimage that contains $X$.
By the same reasoning as in Lemma A.2, $C_{\varepsilon}\subseteq\mathrm{int}\,\,\mathcal{N}_{\delta}(X)$, and hence $C_{\varepsilon}$ is compact. Moreover, by Lemma A.3, every point $x\in\partial C_{\varepsilon}$ satisfies $f(x)=f(X)-\varepsilon$.
Now define $g(x)=\|x-x_{\star}\|^{2}$ and let
$x_{h}\in\mathrm{argmax}_{C_{\varepsilon}}g(x),\quad x_{h}\neq x_{\star},\quad\rho:=\|x_{h}-x_{\star}\|.$
Since the only critical points in $C_{\varepsilon}$ are those in $X$, we have $\nabla f(x_{h})\neq 0$. Thus, there exists an open neighborhood $V$ of $x_{h}$ such that $\Sigma:=\{x\in V:f(x)=f(X)-\varepsilon\}$ is an embedded smooth hypersurface near $x_{h}$. By possibly shrinking $V$, we may assume $V\cap\partial C_{\varepsilon}=\Sigma$. Therefore, maximizing $g(x)$ over $C_{\varepsilon}$ is locally equivalent to maximizing $g(x)$ over the hypersurface $\Sigma$.
Then, by the method of Lagrange multipliers, we obtain
$\nabla g(x_{h})=\lambda\nabla f(x_{h}),$
which implies that there exists $\lambda>0$ such that
$2(x_{\star}-x_{h})=\lambda\nabla f(x_{h}).$
The positivity of $\lambda$ follows from the same reasoning as in Lemma A.4. Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
$x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|},$
which shows that $x_{h}$ is a hallucinated minimizer. ∎
Appendix B Omitted details for Section 3
B.1 Proof of Theorem 3.1
In this subsection, we prove Theorem 3.1. The argument relies on the implicit function theorem, which we state below.
Theorem B.1 (Implicit function theorem, Lee (2013)).
Let $U\subseteq\mathbb{R}^{d}\times\mathbb{R}^{d}$ be an open set, and let $(x,y)$ denote the coordinates on $U$. Suppose $\Phi:U\to\mathbb{R}^{d}$ is a smooth function, $(a,b)\in U$, and $c=\Phi(a,b)$. If the $d\times d$ matrix
$\left(\frac{\partial\Phi^{i}}{\partial y^{j}}(a,b)\right)$
is invertible, then there exist neighborhoods $V_{0}\subseteq\mathbb{R}^{d}$ of $a$ and $W_{0}\subseteq\mathbb{R}^{d}$ of $b$, together with a smooth function $F:V_{0}\to W_{0}$, such that $\Phi^{-1}(c)\cap(V_{0}\times W_{0})$ is the graph of $F$. In other words, $\Phi(x,y)=c$ for $(x,y)\in V_{0}\times W_{0}$ if and only if $y=F(x)$.
Now we prove Theorem 3.1.
See 3.1
Proof.
Let $x_{h}$ be a hallucinated minimizer of $f$ constructed in the proof of Theorem 2.2. Then, it satisfies
$x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{h}+\rho u(x_{h})$
for some perturbation radius $\rho>0$. Let $V$ be an open neighborhood of $x_{h}$ on which the gradient never vanishes. Such a neighborhood exists because $\|\nabla f(x_{h})\|>0$ and $\nabla f$ is continuous. Define $F:\mathbb{R}^{d}\times V\rightarrow\mathbb{R}^{d}$ by
$F(x,y)=y+\rho\ u(y)-x.$
Clearly, $F(x_{\star},x_{h})=0$, and by assumption, $\frac{dF}{dy}(x_{\star},x_{h})=I+\rho\nabla u(x_{h})$ is invertible. By the implicit function theorem, there exist open neighborhoods $U_{0}\subseteq\mathbb{R}^{d}$ of $x_{\star}$ and $V_{0}\subseteq V$ of $x_{h}$, together with a smooth map $G:U_{0}\rightarrow V_{0}$, such that $G(x_{\star})=x_{h}$ and
$F(x,G(x))=0\quad\forall x\in U_{0}.$
This also implies that $G$ is also a local diffeomorphism at $x_{\star}$, since the differential is invertible at $x_{\star}$:
$I+\frac{dF}{dy}\bigg\rvert_{y=x_{h}}\frac{dG}{dx}\bigg\rvert_{x=x_{\star}}=0\iff\frac{dG}{dx}\bigg\rvert_{x=x_{\star}}=\left(\frac{dF}{dy}\right)\bigg\rvert_{y=x_{h}}^{-1}.$
In particular, there exists an open $U\subseteq U_{0}$ around $x_{\star}$ such that $G|_{U}:U\rightarrow G(U)$ is a diffeomorphism. Since $U\cap\mathcal{M}$ is an $m$-dimensional manifold (without boundary), the image $G(U\cap\mathcal{M})$ under the diffeomorphism is also an $m$-dimensional manifold.
Finally, note that any $y\in G(U\cap\mathcal{M})$ satisfies
$x=y+\rho\frac{\nabla f(y)}{\|\nabla f(y)\|},$
where $x=G^{-1}(y)\in\mathcal{M}\cap U$ and hence $x\in\operatorname*{arg\,min}{f}$. Thus, $G(U\cap\mathcal{M})$ forms an $m$-dimensional manifold of hallucinated minimizers. ∎
B.2 Proof of Theorem 3.2 and further discussion
In this subsection, we prove Theorem 3.2 and then discuss the special case of isolated hallucinated minimizers.
See 3.2
Proof.
First, we claim that the set $H$ is closed (hence compact) by construction. Indeed, if $x_{h}\in\bar{H}$, then the corresponding function value must equal $f^{\mathrm{SAM}}(H)=\min f^{\mathrm{SAM}}$. By our assumption on $H$, this implies $x_{h}\in H$. Hence $H=\bar{H}$.
Let $\mathcal{N}_{\delta}(H)$ denote the closed $\delta$-neighborhood of $H$ from the theorem assumption. Since $1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0$ and $\|\nabla f(x_{h})\|>0$ for all $x_{h}\in H$, there exists an open neighborhood $W$ of $H$ such that $1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x)))>0$ and $\|\nabla f(x)\|>0$ for any $x\in W$. By shrinking $\mathcal{N}_{\delta}(H)$ if necessary, we may assume $\mathcal{N}_{\delta}(H)\subseteq W$ and $f^{\mathrm{SAM}}$ is real-analytic on $\mathcal{N}_{\delta}(H)$.
Applying an argument analogous to Lemma A.9, with local maximizers replaced by minimizers, we obtain $\varepsilon^{\star}>0$ such that if $x$ is a critical point in $\mathcal{N}_{\delta}(H)$ with $f^{\mathrm{SAM}}(x)\leq f^{\mathrm{SAM}}(H)+\varepsilon^{\star}$, then $x\in H$.
Now consider the closed neighborhood $\mathcal{N}_{\delta/2}(H)$, and set
$m:=\min_{x\in\partial\mathcal{N}_{\delta/2}(H)}f^{\mathrm{SAM}}(x)>f_{\star},$
where $f_{\star}=f^{\mathrm{SAM}}(H)$. The strict inequality follows from the construction of $\mathcal{N}_{\delta}(H)$. Choose $\varepsilon>0$ such that
$0<\varepsilon<m-f^{\mathrm{SAM}}(H)=m-f_{\star}\quad\textup{ and }\quad 0<\varepsilon<\varepsilon^{\star}.$
Let $C_{\varepsilon}$ be the connected component of the sublevel set
$(f^{\mathrm{SAM}})^{-1}((-\infty,f_{\star}+\varepsilon])=(f^{\mathrm{SAM}})^{-1}([f_{\star},f_{\star}+\varepsilon])$
that contains $H$. Then, $C_{\varepsilon}$ is compact and contains no other critical points of $f^{\mathrm{SAM}}$ besides those in $H$. The proof of $C_{\varepsilon}$ being bounded (hence compact) by $\mathcal{N}_{\delta/2}(H)$ is analogous to Lemma A.2.
Define
$C_{\rho}:=\{x:d(x,C_{\varepsilon})\leq\rho\}.$
Then, $C_{\rho}$ is also compact. Set
$M:=\max_{x\in C_{\rho}}\|\nabla f(x)\|>0,\qquad L:=\max_{x\in\mathcal{N}_{\delta}(H)}\|\nabla^{2}f(x)\|>0,$
and
$\gamma:=\min_{x\in\mathcal{N}_{\delta}(H)}(1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x)))>0.$
Consider the SAM update with fixed $\rho>0$ and constant step size $\eta_{k}=\eta$ chosen such that
$0<\eta<\min\Big\{\frac{\delta}{2M},\ \frac{2\gamma}{L}\Big\},$
and initialization at $x_{0}\in C_{\varepsilon}$. We show by induction that the SAM iterates $\{x_{k}\}$ remain in $C_{\varepsilon}$. The base case $x_{0}\in C_{\varepsilon}$ is true by assumption. Suppose $x_{k}\in C_{\varepsilon}$. Then, by definition of the SAM update, $x_{k}^{+}\in C_{\rho}$ and $x_{k+1}=x_{k}-\eta\nabla f(x_{k}^{+})$. We claim $x_{k+1}\in\mathcal{N}_{\delta}(H)$, since
$d(x_{k+1},H)$ $=\inf_{x_{h}\in H}\|x_{k+1}-x_{h}\|$
$\leq\inf_{x_{h}\in H}\|x_{k}-x_{h}\|+\|x_{k+1}-x_{k}\|$
$\leq\frac{\delta}{2}+\eta\|\nabla f(x_{k}^{+})\|$
$\leq\frac{\delta}{2}+\frac{\delta}{2M}\cdot M$
$\leq\delta.$
By LL–smoothness of $f^{\mathrm{SAM}}$ on $\mathcal{N}_{\delta}(H)$,
$f^{\mathrm{SAM}}(x_{k+1})$ $\leq f^{\mathrm{SAM}}(x_{k})+\langle\nabla f^{\mathrm{SAM}}(x_{k}),\,x_{k+1}-x_{k}\rangle+\frac{L}{2}\|x_{k+1}-x_{k}\|^{2}$
$=f^{\mathrm{SAM}}(x_{k})-\eta\langle\nabla f^{\mathrm{SAM}}(x_{k}),\nabla f(x_{k}^{+})\rangle+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}.$
Since
$\nabla f^{\mathrm{SAM}}(x)=(I+\rho\nabla u(x))\nabla f(x^{+}),$
we obtain
$f^{\mathrm{SAM}}(x_{k+1})$ $\leq f^{\mathrm{SAM}}(x_{k})-\eta\langle\nabla f^{\mathrm{SAM}}(x_{k}),\nabla f(x_{k}^{+})\rangle+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}$
$=f^{\mathrm{SAM}}(x_{k})-\eta\rho\langle\nabla u(x_{k})\nabla f(x_{k}^{+}),\nabla f(x_{k}^{+})\rangle-\eta\|\nabla f(x_{k}^{+})\|^{2}+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}$
$\leq f^{\mathrm{SAM}}(x_{k})-\eta\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{k})))\|\nabla f(x_{k}^{+})\|^{2}-\eta\|\nabla f(x_{k}^{+})\|^{2}+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}$
$\leq f^{\mathrm{SAM}}(x_{k})-\eta\left(1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{k})))-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}$
$\leq f^{\mathrm{SAM}}(x_{k})-\eta\left(\gamma-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}$
$<f^{\mathrm{SAM}}(x_{k}),$
where the last inequality follows from $0<\eta<\frac{2\gamma}{L}$. Moreover, since the descent property $f^{\mathrm{SAM}}(x_{k+1})<f^{\mathrm{SAM}}(x_{k})$ holds when $\eta$ is replace by $\eta t$ for $t\in[0,1]$, the line segment from $x_{k}$ to $x_{k+1}$ lies within the sublevel set $(f^{\mathrm{SAM}})^{-1}([f_{\star},f_{\star}+\varepsilon])$. Thus, because $x_{k}\in C_{\varepsilon}$, we conclude $x_{k+1}\in C_{\varepsilon}$ by the connectedness of $C_{\varepsilon}$. This completes the induction.
Finally, since $f^{SAM}(x_{k})$ is decreasing and bounded below, we have $\eta\left(\gamma-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|\rightarrow 0$. Hence, $\nabla f(x_{k}^{+})\to 0$. If $x_{\infty}$ is a limit point of $\{x_{k}\}_{k=0,1,\dots}$, then $\nabla f(x_{\infty}^{+})=0$. By construction of $C_{\varepsilon}$, this implies $x_{\infty}\in H$. Therefore, $d(x_{k},H)\to 0$. ∎
To discuss point convergence to isolated hallucinated minimizers, we now turn to the case where the manifold $\mathcal{M}$ in Theorem 3.1 reduces to a single isolated point, i.e., 0-dimensional. In this case, we can show that the corresponding hallucinated minimizer is also isolated. This can be viewed as the special case where $H$ in Theorem 3.2 is a singleton.
Lemma B.2.
Let $x_{\star}$ be a minimizer of a $C^{1}$ function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying the assumptions of Theorem 2.1. Suppose $x_{\star}$ is an isolated minimizer; that is, there exists an open neighborhood of $x_{\star}$ in which it is the unique critical point and the unique minimizer. Let $x_{h}$ be a hallucinated minimizer of $f$ constructed in the proof of Theorem 2.1 with $\rho>0$. If $I+\rho\nabla u(x_{h})$ is invertible, then there exists an open neighborhood $W$ of $x_{h}$ such that
  • $\nabla f$ never vanishes on $W$, and thus $f^{\mathrm{SAM}}$ is well-defined on $W$;
  • no point other than $x_{h}$ satisfies $\nabla f(x^{+})=0$ in $W$; and
  • $x_{h}$ is the unique hallucinated minimizer in $W$.
Such a point $x_{h}$ is called an isolated hallucinated minimizer.
Proof.
Since $x_{h}$ is a hallucinated minimizer of $f$ constructed in the proof of Theorem 2.1, it satisfies
$x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{h}+\rho u(x_{h})$
for some perturbation radius $\rho>0$. Let $U_{0}$ and $V_{0}$ be open neighborhoods of $x_{\star}$ and $x_{h}$, respectively, as constructed in the proof of Theorem 3.1. Also, let $U_{1}$ be an open neighborhood of $x_{\star}$ that contains no other minimizers, by assumption. Define $W:=G(U_{0}\cap U_{1})$, where $G$ is the $C^{1}$ mapping constructed in the proof of Theorem 3.1. We claim that $W$ is the desired open neighborhood of $x_{h}$.
First, $W$ is open since $G$ is a local diffeomorphism at $x_{\star}$. Moreover, $f^{\mathrm{SAM}}$ is well-defined on $W$ by the construction of $V_{0}$. Suppose $y\in W$ satisfies $\nabla f(y^{+})=0$. Then, since
$x=y+\rho\frac{\nabla f(y)}{\|\nabla f(y)\|}$
for the unique $x=G^{-1}(y)\in U_{1}$, it follows that $\nabla f(y^{+})=\nabla f(x)=0$, contradicting the fact that $x$ is an isolated minimizer. Similarly, if $y$ is a hallucinated minimizer, then the uniqueness of $x=G^{-1}(y)$ implies $x=x_{\star}$, and hence $y=x_{h}$. This proves the claim. ∎
Then, as a corollary of Theorem 3.2 and the previous lemma, we obtain the following point convergence result.
Corollary B.3.
Assume $f$ satisfies the assumptions in Theorem 2.1 and, in addition, $f\in C^{2}$. Let $x_{h}$ be an isolated hallucinated minimizer of $f$ for a fixed perturbation radius $\rho>0$, constructed in the proof of Theorem 2.1. Suppose
$1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0.$
Then, the SAM iterates, when initialized sufficiently close to $x_{h}$, converge to $x_{h}$ for sufficiently small fixed step size $\eta_{k}=\eta$.
The real-analytic property of $f$ is used to apply the Łojasiewicz inequality in order to construct a neighborhood where critical points do not accumulate. However, since Lemma B.2 guarantees the existence of isolated hallucinated minimizers without this assumption, the real-analytic condition is not required here. Hence, the $C^{2}$ assumption on $f$ is sufficient to ensure the existence of constants $L$ and $\gamma$ as in the proof of Theorem 3.2.
Appendix C Experimental details for SAM in deep learning
C.1 SAM with full-batch gradients
In Section 4.1, we train a neural network using SAM with full-batch gradients. Specifically, the model is a two-layer network with 128 hidden units and Tanh activations, trained on the MNIST dataset (LeCun et al., 1998). The classification task uses cross-entropy loss. Training is implemented in PyTorch (Paszke et al., 2019) with a learning rate of $0.01$, momentum $0.9$, and no weight decay. We run $20$ million updates with perturbation radius $\rho=1.8$ to obtain the convergence point $x_{h}$, whose loss landscape is shown in Figure 4. In this subsection, we provide additional visualizations to further examine the local properties of $x_{h}$.
Figure 7(a) presents a one-dimensional view along the line connecting $x_{h}$ and $x_{h}^{+}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$, parameterized as $x(\alpha)=(1-\alpha)x_{h}+\alpha x_{h}^{+}$. The plot shows that $x_{h}$ is not a minimizer of the original objective and that the surrounding loss landscape differs substantially from that around $x_{h}^{+}$. This demonstrates that the phenomenon of SAM converging to a hallucinated minimizer is fundamentally distinct from the case in which a saddle point becomes an attractor, which requires the surrounding quadratic structure to hold (Compagnoni et al., 2023).
(a) Original loss $f$
(b) SAM loss $f^{\mathrm{SAM}}$
Figure 7: Visualizations of the hallucinated minimizer $x_{h}$: (a) original loss $f$ along the line between $x_{h}$ and $x_{h}^{+}$; (b) SAM loss $f^{\mathrm{SAM}}$ over the plane defined by $x_{h}$, $x_{N}$, and $x_{N}^{\prime}$.
Figure 7(b) extends Figure 4(c) with an additional visualization of the SAM loss. We initialize $x_{0}$ by adding a small random perturbation of magnitude $0.1$ to $x_{h}$, and then perform $N=1000$ SAM steps, yielding the same $x_{N}$ reported in Figure 4(c). Applying an independent perturbation followed by the same procedure gives $x_{N}^{\prime}$. We then consider the plane spanned by $x_{h}$, $x_{N}$, and $x_{N}^{\prime}$, parameterized as $x(\alpha,\beta)=x_{h}+\alpha u+\beta v$ with $u=x_{N}-x_{h}$ and $v$ chosen orthogonal to $u$. On this plane, the visualization shows that the hallucinated minimizers are not confined to a one-dimensional curve but instead extend into a two-dimensional surface-like structure.
(a) Final train loss
(b) Final test accuracy
Figure 8: Comparison of SAM-only and the switching strategy across different perturbation radii $\rho$. Results are obtained using SAM with full-batch gradients over 80 seeds. Bold lines indicate the mean, and shaded areas represent the standard deviation.
In the experiments reported in Figure 5, we investigate SAM with full-batch gradients by varying both the perturbation radius and the random seeds. Under the same experimental setting, Figure 8 shows the final training loss and test accuracy at the last step for perturbation radii $\rho=1.0,1.1,\ldots,2.0$, evaluated across 80 seeds. The results demonstrate that the performance of SAM is highly sensitive to the perturbation radius, whereas the switching strategy maintains stable performance even for larger values of $\rho$.
$\rho=0.1$
$\rho=0.4$
$\rho=0.7$
$\rho=1.0$
(a) SAM-only
(b) SGD $\rightarrow$ SAM
Figure 9: Training loss and test accuracy curves for SAM with stochastic gradients on CIFAR-100 using ResNet-18. The top row shows SAM-only training, while the bottom row applies a switching strategy that runs stochastic gradient descent for the first 10% of epochs before switching to SAM. Columns correspond to perturbation radii $\rho\in\{0.1,0.4,0.7,1.0\}$. Final average test accuracies are 74.29%, 76.36%, 69.64%, and 61.58% for SAM-only, compared to 74.48%, 77.22%, 73.35%, and 67.93% for the switching strategy.
C.2 SAM with stochastic gradients
We examine whether the phenomena observed with full-batch SAM also arise in the stochastic setting, as shown in Figure 6. ResNet-18 is trained on CIFAR-100 with standard data augmentations, including random cropping with padding, horizontal flipping, and Cutout (DeVries & Taylor, 2017). The mini-batch size is $64$, the learning rate $0.01$, momentum $0.9$, and weight decay $10^{-4}$. Training proceeds for $200$ epochs with cosine-annealed learning rates (Loshchilov & Hutter, 2016), following the practical implementation of FSAM (Li et al., 2024). The switching strategy applies plain stochastic gradient descent for the first $10\%$ of epochs before switching to stochastic SAM.
Under the same setting, Figure 9 reports the CIFAR-100 results, comparing SAM-only with the switching strategy. Experiments are conducted for perturbation radii $\rho=0.1,0.4,0.7,1.0$ across $16$ random seeds. Each curve shows training loss and test accuracy over epochs, with bold lines denoting the mean across seeds and shaded regions indicating the standard deviation. The results show that, as in the full-batch case, SAM performance degrades with larger perturbation radii, whereas the switching strategy remains stable and robust across all settings.
Appendix D Two-dimensional synthetic function for visualization
To visualize how the SAM perturbation radius $\rho$ affects the objective, we introduce the following two-dimensional synthetic function (originally illustrated in Figure 1):
$f(x,y)=0.8\exp\left(-\frac{x^{2}+y^{2}}{(2.5)^{2}}\right)\cdot W_{X}(x)-\exp\left(-(x+1.55\cos(y/1.5))^{2}\right)\cdot W_{Y}(y)+1,$
where
$W_{X}(x)=\begin{cases}0,&x\leq-1,\\ 0.5\left(1-\cos(\pi(x+1))\right),&-1<x<0,\\ 1,&x\geq 0,\end{cases}$
$W_{Y}(y)=\begin{cases}1,&|y|\leq 0.6,\\ 0.5\left(1+\cos\left(\pi\cdot\frac{|y|-0.6}{5.0}\right)\right),&0.6<|y|<5.6,\\ 0,&|y|\geq 5.6.\end{cases}$
The function $f(x,y)$ is continuously differentiable, since all its components are smoothly joined. It is designed so that its minimizer set forms a curve. In fact, the global minimizer set of $f$ is exactly
$\left\{(x,y)\in\mathbb{R}^{2}\;\middle|\;x=-1.55\cos\left(\tfrac{y}{1.5}\right),\;|y|\leq 0.6\right\}.$
Figure 1 shows the original function $f$, the SAM objective $f^{\mathrm{SAM}}$, and its gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ at perturbation radius $\rho=2.8$. To further examine the effect of the perturbation radius, Figures 10 and 11 illustrate how the SAM objective $f^{\mathrm{SAM}}$ and its gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ evolve as $\rho$ varies over the range $0,0.5,\ldots,3.5$. In this setting, the SAM minimizers are defined as the regions where $f^{\mathrm{SAM}}$ attains its minimum values; in practice, they appear either as isolated points or as continuous curve-like structures.
(a) $\rho=0.0$
(b) $\rho=0.5$
(c) $\rho=1.0$
(d) $\rho=1.5$
(e) $\rho=2.0$
(f) $\rho=2.5$
(g) $\rho=3.0$
(h) $\rho=3.5$
Figure 10: SAM objective $f^{\mathrm{SAM}}(x)$ under different perturbation radii $\rho$. The corresponding SAM minimizers are shown in pink.
(a) $\rho=0.0$
(b) $\rho=0.5$
(c) $\rho=1.0$
(d) $\rho=1.5$
(e) $\rho=2.0$
(f) $\rho=2.5$
(g) $\rho=3.0$
(h) $\rho=3.5$
Figure 11: SAM gradient field $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ for different perturbation radii $\rho$. The corresponding SAM minimizers are shown in pink.
An analysis of the SAM minimizers as a function of the perturbation radius $\rho$ reveals two distinct regimes. For small $\rho$, the minimizers approach the critical points of $f$. Although $f^{\mathrm{SAM}}$ is not defined at a critical point, higher-resolution numerical experiments show convergence arbitrarily close to such points. At $\rho=0$, $f^{\mathrm{SAM}}$ coincides with $f$, and the minimizers exactly match those of $f$. For $\rho=0.5$, the minimizers remain on the original minimizer set, whereas for $\rho=1.0$ and $\rho=1.5$, they shift toward the maximizers of $f$. The corresponding gradient fields indicate that these critical points act as attractors of the SAM dynamics, consistent with the theoretical analysis of Compagnoni et al. (2023).
In contrast, for larger perturbation radii, the SAM minimizers form a curve on the right-hand side, starting near the maximizer and drifting outward as $\rho$ increases. This behavior is consistent with the proof of Theorem 2.1, which shows that hallucinated minimizers emerge for a sufficiently large perturbation radius $\rho$. Theorem 3.1 further establishes that these minimizers preserve the dimensionality of the original minimizer manifold. Meanwhile, the SAM gradient field shows that these minimizers act as attractors within their neighborhood, and the conditions of Theorem 3.2 are indeed satisfied. Taken together, these observations show that our theory offers a full theoretical explanation of the empirical phenomena of hallucinated minimizers in this example.

### v2

@@
 Abstract
-Sharpness-Aware Minimization (SAM) is a widely used method that steers training toward flatter minimizers, which typically generalize better.
-In this work, however, we show that SAM can converge to hallucinated minimizers—points that are not minimizers of the original objective.
-We theoretically prove the existence of such hallucinated minimizers and establish conditions for local convergence to them.
-We further provide empirical evidence demonstrating that SAM can indeed converge to these points in practice.
-Finally, we propose a simple yet effective remedy for avoiding hallucinated minimizers.
+Sharpness-Aware Minimization (SAM) is widely used to seek flatter minima—often linked to better generalization.
+In its standard implementation, SAM updates the current iterate using the loss gradient evaluated at a point perturbed by distance $\rho$ along the normalized gradient direction.
+We show that, for some choices of $\rho$, SAM can stall at points where this shifted (perturbed-point) gradient vanishes despite a nonzero original gradient, and therefore, they are not stationary points of the original loss.
+We call these points hallucinated minimizers, prove their existence under simple nonconvex landscape conditions (e.g., the presence of a local minimizer and a local maximizer), and establish sufficient conditions for local convergence of the SAM iterates to them.
+We corroborate this failure mode in neural network training and observe that it aligns with SAM’s performance degradation often seen at large $\rho$.
+Finally, as a practical safeguard, we find that a short initial SGD warm-start before enabling SAM mitigates this failure mode and reduces sensitivity to the choice of $\rho$.
 (a) Landscape of $f$
@@
 (a) Smooth function $f$ with a minimizer set and an isolated maximizer.
-(b) $f^{\text{SAM}}=f(x+\rho\tfrac{\nabla f(x)}{\|\nabla f(x)\|})$; its minimizers do not correspond to minimizers or stationary points of $f$ and are therefore hallucinated.
-(c) Vector field of $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$; the hallucinated minimizers are attractors of the SAM iteration.
+(b) $f^{\mathrm{SAM}}=f(x+\rho\,u(x))$; its minimizers (labeled as “SAM minimizers” in the plot) do not correspond to minimizers or stationary points of $f$ and are therefore hallucinated.
+(c) Vector field of the SAM gradient $\nabla f(x^{+})$; the hallucinated minimizers are attractors of the SAM iteration.
 1 Introduction
-It has been empirically observed in deep learning that flat minimizers tend to generalize better than sharp ones (Neyshabur et al., 2017; Jiang et al., 2020).
-Motivated by this observation, Sharpness-Aware Minimization (SAM) was proposed as a training method that fits the data while simultaneously regularizing against the sharpness (Foret et al., 2021; Zheng et al., 2021; Wu et al., 2020).
-Specifically, for a differentiable function $f\colon\mathbb{R}^{d}\to\mathbb{R}$, SAM minimizes the objective
-$f^{\mathrm{SAM}}(x):=f\!\left(x+\rho\,\frac{\nabla f(x)}{\|\nabla f(x)\|}\right),$
-where the perturbation radius $\rho>0$ controls the strength of the regularization.
-By encouraging convergence to flat minimizers, SAM has demonstrated strong empirical performance and has inspired a variety of practical variants (Foret et al., 2021).
-However, although prior theoretical studies have analyzed the convergence of SAM to a minimizer under various conditions, most rely on assumptions that rarely hold in deep learning, such as convexity (Si & Yun, 2024) or a decaying perturbation radius (Khanh et al., 2024; Oikonomou & Loizou, 2025).
-Consequently, the validity of these convergence results is not guaranteed in practical settings.
+Sharpness-Aware Minimization (SAM) is widely used to bias training toward flatter regions of the loss landscape (Foret et al., 2021), often correlated with improved generalization (Neyshabur et al., 2017; Jiang et al., 2020).
+For a differentiable objective $f:\mathbb{R}^{d}\to\mathbb{R}$, SAM is commonly motivated by the min–max formulation
+$\min_{x}\max_{\|\epsilon\|\leq\rho}f(x+\epsilon),$
+and the inner maximization is typically approximated by a first-order perturbation along the normalized gradient direction. This leads to the surrogate (“SAM”) loss
+$f^{\mathrm{SAM}}(x):=f\!\left(x+\rho\,u(x)\right),\qquad u(x)=\frac{\nabla f(x)}{\|\nabla f(x)\|},$
+where $\rho>0$ is the perturbation radius controlling the strength of the sharpness regularization.
+A practical mismatch in SAM updates.
+A central subtlety—often glossed over in high-level descriptions of SAM—is that the standard implementation does not follow the surrogate gradient $\nabla f^{\mathrm{SAM}}(x)$.
+Instead, it updates using the shifted gradient evaluated at the perturbed point:
+$x_{k+1}=x_{k}-\eta_{k}\,\nabla f(x_{k}^{+}),\qquad x^{+}:=x+\rho\,u(x).$
+Consequently, the practical SAM dynamics can become stationary whenever $\nabla f(x^{+})=0$, even if $\nabla f(x)\neq 0$.
+This mismatch becomes especially relevant in the large-$\rho$ regime: in practice, $\rho$ is tuned by sweeps or grid search, and prior work reports performance degradation and instability at larger radii on standard benchmarks (Andriushchenko and Flammarion, 2022).
+On the theory side, most convergence guarantees assume $\rho$ is below a threshold under various regularity conditions, leaving the large-$\rho$ regime comparatively less understood (Si and Yun, 2023; Khanh et al., 2024; Oikonomou and Loizou, 2025).
+Hallucinated minimizers.
+We show that the shifted-gradient dynamics can admit a qualitatively distinct failure mode: SAM can stall at points where $\nabla f(x^{+})=0$ despite $\nabla f(x)\neq 0$.
+We call such points hallucinated minimizers if, in addition, they are local minimizers of the surrogate loss $f^{\mathrm{SAM}}$.
+Since $\nabla f(x)\neq 0$, hallucinated minimizers are not stationary points of the original objective $f$ (Figure 1).
+This failure mode is inherently nonconvex (it cannot occur for convex $f$), and becomes particularly relevant in the large-$\rho$ regime, where the empirical performance of SAM is often observed to degrade (Andriushchenko and Flammarion, 2022).
+Our results.
+We provide a geometric and dynamical characterization of hallucinated minimizers.
+Specifically, we prove that hallucinated minimizers exist for a nontrivial interval of radii, including values in a large-$\rho$ regime (Theorem 2.2) under simple nonconvex landscape conditions (e.g., the coexistence of a local minimizer and a local maximizer).
+We extend this existence result from isolated maximizers to bounded local maximizer sets under a real-analyticity assumption (Theorem 2.4).
+We further show that hallucinated minimizers are not merely an algebraic possibility: an isolated connected component can be locally attracting for the discrete-time SAM dynamics (Theorem 3.1).
+Finally, we show that hallucinated minimizers can be non-isolated and inherit the dimension of minimizer manifolds via local invertibility of the perturbation map (Theorem 3.2).
+Empirically, we provide evidence consistent with this failure mode on neural network objectives using diagnostics that compare shifted vs.
+unshifted gradients and losses: in smooth full-batch settings we directly verify convergence to points with nearly vanishing shifted gradient but nonzero original gradient, and we observe similar mismatch signatures in larger networks beyond the smooth/analytic setting.
+Finally, as a practical safeguard, we report that a short initial SGD warm-start before enabling SAM mitigates this failure mode and reduces sensitivity to the choice of $\rho$.
+Our formal results concern deterministic (full-batch) dynamics; stochastic experiments are included only as empirical observations.
 Contributions.
-In this paper, we theoretically and empirically demonstrate that SAM can, in fact, hallucinate minimizers, in the sense that its iterates may converge to points that are not minimizers of the original objective, as illustrated in Figure 1.
-This finding reveals a previously unrecognized failure mode of SAM in deep learning, one that is fundamentally distinct from issues caused by local minima or saddle points.
-To address this, we present a simple switching strategy that effectively avoids hallucinated minimizers.
+  • We identify and formalize hallucinated minimizers: points where the shifted gradient vanishes, $\nabla f(x^{+})=0$, even though the original gradient is nonzero, $\nabla f(x)\neq 0$, and $x$ is a local minimizer of $f^{\mathrm{SAM}}(x)=f(x^{+})$.
+  • We prove existence of hallucinated minimizers for a nontrivial interval of radii (including a large-$\rho$ regime) under mild nonconvex landscape conditions, and extend the result to bounded local maximizer sets under real-analyticity.
+  • We provide sufficient conditions under which an isolated hallucinated-minimizer component is locally attracting for the discrete-time SAM dynamics.
+  • We show that hallucinated minimizers can form smooth manifolds inheriting the dimension of minimizer manifolds via local invertibility of the perturbation map.
+  • We provide empirical diagnostics and observations in neural network training that are consistent with the proposed mechanism.
 1.1 Related work
-SAM and its variants.
-Building on the observation that flat minimizers are stable under small perturbations (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017; Chaudhari et al., 2019), SAM was proposed as a method to seek such minimizers (Foret et al., 2021).
-SAM has demonstrated remarkable performance across a wide range of deep learning tasks (Foret et al., 2021; Bahri et al., 2022; Zhong et al., 2022; Lee et al., 2023), motivating numerous extensions.
-One line of work improves SAM by modifying its perturbation direction.
-Kwon et al.
-(2021) adapted the perturbation in a scale-invariant manner, while Kim et al.
-(2022) redefined it using Fisher information geometry.
-Li et al.
-(2024) removed the full-batch gradient from the perturbation direction and leveraged stochastic gradient noise for generalization.
-Li & Giannakis (2024) incorporated momentum into the perturbation to suppress variance and stabilize the adversary.
-Instead of altering the perturbation itself, Zhuang et al.
-(2022) adjusted the gradient update through orthogonal decomposition to reduce the surrogate gap.
-Another line of research addresses the computational overhead of SAM, which stems from requiring two gradient computations per step.
-Some approaches reduce the number of gradient evaluations: Liu et al.
-(2022) computed the perturbation only periodically, Jiang et al.
-(2023) activated SAM only when the gradient norm is large, and Du et al.
-(2022b) reused gradients to avoid the second computation.
-Others restrict the scope of perturbations: Du et al.
-(2022a) applied them only to a random subset of parameters and sharpness-sensitive data, while Mueller et al.
-(2023) showed that limiting them to normalization layers preserves most of the benefits.
-Beyond reducing computation, Xie et al.
-(2024) improved training efficiency by parallelizing the two gradient computations.
+SAM and practical variants.
+SAM was introduced as a method to regularize sharpness by adversarially perturbing parameters during training (Foret et al., 2021).
+A large body of follow-up work proposes variants that modify the perturbation geometry or the update rule, e.g., scale-invariant perturbations (ASAM) (Kwon et al., 2021), Fisher-geometry perturbations (Kim et al., 2022), incorporation of momentum into the perturbation (Li and Giannakis, 2023) or modified updates such as orthogonal decomposition (GSAM) (Zhuang et al., 2022).
+Other works reduce SAM’s two-backward-pass overhead via approximation, partial perturbations, or scheduling (e.g., (Liu et al., 2022; Jiang et al., 2023; Du et al., 2022a, b; Mueller et al., 2023)).
+In this paper, we focus on the standard SAM update—the shifted-gradient step used in common implementations—and analyze its behavior in the large-$\rho$ regime.
 Theoretical analyses of SAM.
-Alongside its practical success, a growing body of theoretical work has analyzed SAM and examined its generalization properties from multiple perspectives.
-Wen et al.
-(2022) formalized the precise notion of sharpness minimized by SAM, clarifying its regularization effect.
-Möllenhoff & Khan (2023) reinterpreted SAM as a relaxation of Bayesian inference.
-Chen et al.
-(2023) showed that SAM mitigates noise fitting and improves generalization over stochastic gradient descent.
-Wei et al.
-(2023) further demonstrated that SAM alone can enhance adversarial robustness while maintaining clean accuracy.
-Another line of work investigates the training dynamics and stability of SAM.
-Compagnoni et al.
-(2023) analyzed SAM through the lens of stochastic differential equations, offering a continuous-time perspective.
-Bartlett et al.
-(2023) studied quadratic objectives, showing how SAM oscillates across narrow valleys before drifting toward wider minimizers.
-Dai et al.
-(2023) demonstrated that normalization plays a key role in stabilizing SAM and ensuring robustness.
-Long & Bartlett (2024) extended the edge-of-stability threshold of gradient descent to SAM, showing that it depends on the gradient norm.
-More recently, Zhou et al.
-(2025) highlighted a late-phase effect, whereby SAM selects flatter minimizers when applied in the later stages of training.
-Finally, several works have investigated the convergence properties of SAM in diverse settings.
-Andriushchenko & Flammarion (2022) proposed USAM, an unnormalized variant obtained by removing gradient normalization, and analyzed its convergence.
-Si & Yun (2024) provided a systematic study across convex, strongly convex, and nonconvex regimes.
-Khanh et al.
-(2024) developed a convergence analysis of SAM and its variants within the framework of inexact gradient descent.
-Most recently, Oikonomou & Loizou (2025) analyzed SAM and USAM within a unified framework and proved convergence under the Polyak–Łojasiewicz condition.
-Hallucinated minimizers and our contribution.
-Several prior studies have reported that SAM may converge to points that are not minimizers of the original loss.
-Kaddour et al.
-(2022) empirically observed that SAM can become trapped at saddle points, and Kim et al.
-(2023); Compagnoni et al.
-(2023) provided a theoretical explanation of this phenomenon using a continuous-time model of SAM.
-In another line of work, Bartlett et al.
-(2023) showed that the SAM update is equivalent to gradient descent on a surrogate function in the quadratic case, and Si & Yun (2024) proposed a virtual loss to extend this idea, although it is rigorously defined only in one dimension and lacks guarantees in higher dimensions.
-In contrast, hallucinated minimizers represent a fundamentally different failure mode of SAM.
-They differ from saddle points in that they are not critical points of the original loss, and from surrogate-based interpretations in that they arise directly from the SAM objective.
-Our analysis applies to general nonconvex and high-dimensional settings and provides a rigorous characterization of hallucinated minimizers that can emerge in practical deep learning scenarios.
-1.2 Preliminaries and notation
-Throughout the paper, we denote $u(x):=\nabla f(x)/\|\nabla f(x)\|$ (for $x$ such that $\nabla f(x)\neq 0$).
-To optimize the SAM objective $f^{\mathrm{SAM}}(x)$, we require its gradient $\nabla f^{\mathrm{SAM}}(x)$.
-Under the standard smoothness assumption on $f$, this gradient is given by
-$\nabla f^{\mathrm{SAM}}(x)=\big(I+\rho\,\nabla u(x)\big)\,\nabla f\big(x+\rho\,u(x)\big),$
-where $\nabla u(x)$ denotes the Jacobian of $u(x)$.
-In practice, however, for computational simplicity, one does not use the exact gradient $\nabla f^{\mathrm{SAM}}$.
-Instead, SAM employs the “shifted” gradient $\nabla f\big(x+\rho\,u(x)\big)$.
-This yields the (full-batch) SAM iteration
-$x_{k+1}=x_{k}-\eta_{k}\,\nabla f\big(x_{k}+\rho\,u(x_{k})\big),\qquad k=0,1,2,\dots,$
-where $x_{0}\in\mathbb{R}^{d}$ is the starting point and $\eta_{0},\eta_{1},\ldots\in\mathbb{R}_{+}$ is the sequence of step sizes.
-The perturbation radius $\rho>0$ controls the degree of “flatness”: larger values of $\rho$ expand the neighborhood over which the loss is minimized, thereby steering the SAM iteration towards flatter minimizers.
-In practice, SAM is typically implemented with stochastic gradients.
-When clarification is needed, we refer to SAM with full-batch gradients versus SAM with stochastic gradients.
-We introduce some notation.
+Existing theory has examined SAM from multiple perspectives, including sharpness notions and generalization interpretations (Wen et al., 2023; Möllenhoff and Khan, 2023; Chen et al., 2023; Wei et al., 2023).
+Another line studies SAM’s dynamics and stability, e.g., via continuous-time and SDE approximations (Compagnoni et al., 2023), quadratic dynamics and surrogate interpretations (Bartlett et al., 2023; Si and Yun, 2023), and stability phenomena related to normalization and edge-of-stability effects (Dai et al., 2023; Long and Bartlett, 2024).
+Several works provide convergence guarantees for SAM and related methods under various assumptions, typically requiring a small-radius condition or a regularity condition that effectively bounds the perturbation effect (e.g., convex/strongly convex regimes, inexact gradient descent frameworks, or PL-type conditions) (Andriushchenko and Flammarion, 2022; Si and Yun, 2023; Khanh et al., 2024; Oikonomou and Loizou, 2025).
+In contrast, our results characterize a distinct phenomenon that can arise at large $\rho$: the standard SAM dynamics can stall at points that are not stationary (and hence not locally optimal) for $f$.
+Differences from prior work: hallucinated minimizers vs. saddle trapping.
+Prior work has reported that SAM can become trapped near saddle points (Kaddour et al., 2022), and theoretical explanations have been developed via continuous-time SAM models (Kim et al., 2023; Compagnoni et al., 2023).
+Hallucinated minimizers represent a qualitatively different failure mode: saddle trapping corresponds to $\nabla f(x)=0$ (stationarity of the original loss), whereas hallucinated minimizers satisfy $\nabla f(x)\neq 0$ but $\nabla f(x^{+})=0$, so they are stationary only for the SAM’s shifted-gradient update.
+Moreover, our analysis provides $(i)$ geometric existence mechanisms tied to nonconvex landscapes and large $\rho$, $(ii)$ sufficient conditions for local attraction under the discrete-time SAM dynamics, and $(iii)$ manifold structure results via local invertibility of the perturbation map.
+Switching schedules.
+Switching schedules for SAM have been explored in prior work, often motivated by computational overhead or late-phase effects (Andriushchenko and Flammarion, 2022; Zhou et al., 2025).
+Complementary to these perspectives, we report that a short early SGD warm-start before enabling SAM can mitigate the hallucinated-minimizer failure mode identified in this paper.
+Our goal is not to propose a new training recipe, but to provide a simple and broadly compatible safeguard aligned with our mechanism.
+1.2 Notation and terminology
+We use $\|\cdot\|$ for the Euclidean norm (and its induced operator norm for matrices), write $\mathrm{Sym}(A):=(A+A^{\top})/2$, and let $\lambda_{\min}(A)$ denote the smallest eigenvalue of a symmetric matrix $A$.
 A function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is called real-analytic if its Taylor series at any point $x_{0}$ converges to $f$ on a neighborhood of $x_{0}$.
 For $\alpha\in\mathbb{R}$, the $\alpha$-superlevel set of $f$ is defined as $\{x:\,f(x)\geq\alpha\}$.
-For a set $C\subset\mathbb{R}^{d}$, we write $\partial C$ for its boundary.
-A set $C$ is connected if it cannot be expressed as the union of two disjoint, nonempty open sets.
+For a set $C\subset\mathbb{R}^{d}$, we write $\partial C$ for its boundary, and we say that $C$ is connected if it cannot be expressed as the union of two disjoint, nonempty open sets.
 The distance from a point $x\in\mathbb{R}^{d}$ to a nonempty set $C\subseteq\mathbb{R}^{d}$ is $d(x,C):=\inf_{y\in C}\|x-y\|$; if $C$ is closed, the infimum is attained, and hence $d(x,C)=\min_{y\in C}\|x-y\|$.
-For $\delta>0$, $B_{\delta}(x)$ is the ball centered at $x$ with radius $\delta$ and the (closed) $\delta$-neighborhood of $C$ is $\mathcal{N}_{\delta}(C)=\{x:\,d(x,C)\leq\delta\}$.
+For $\delta>0$, $B_{\delta}(x)$ is the open ball centered at $x$ with radius $\delta$ and $\mathcal{N}_{\delta}(C)=\{x:\,d(x,C)\leq\delta\}$ is the (closed) $\delta$-neighborhood of $C$.
 2 Existence of hallucinated minimizers
-In this section, we establish—under very mild assumptions—the existence of hallucinated minimizers: local minimizers of $f^{\mathrm{SAM}}$ that are not even stationary points of the original function $f$.
-Formally, we define hallucinated minimizers as follows:
-Definition.
-A point $x\in\mathbb{R}^{d}$ is a hallucinated minimizer of $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ for $\rho>0$ if $x$ is a local minimizer of $f^{\mathrm{SAM}}=f\Big(x+\rho\tfrac{\nabla f(x)}{\|\nabla f(x)\|}\Big)$ while satisfying $\nabla f(x)\neq 0$.
-When the loss function $f$ is convex, hallucinated minimizers cannot arise, as shown in Theorem A.1 of Appendix A.
-In practical deep learning, however, the loss function $f$ is highly nonconvex, and local maximizers of $f$ can give rise to hallucinated minimizers.
-2.1 Simplified existence proof with isolated maximizers
-We begin with a simplified proof under the more restrictive assumption that $f$ has an isolated local maximizer, defined as a point $x^{\bullet}$ with an open neighborhood $U$ such that
-$\nabla f(x)\neq 0\,\,\text{ and }\,\,f(x)<f(x^{\bullet})\quad\text{for all }x\in U\setminus\{x^{\bullet}\}.$
-Theorem 2.1.
-Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be continuously differentiable.
-Assume $f$ has a global minimizer (not necessarily unique) and an isolated local maximizer.
-Then, a hallucinated minimizer exists for some $\rho>0$.
+In this section, under simple nonconvex landscape assumptions, we establish the existence of hallucinated minimizers: points where the (full-batch) standard SAM update can become stationary even though the original gradient is nonzero.
+Why a mismatch is possible.
+Recall that SAM is commonly motivated by the surrogate objective
+$f^{\mathrm{SAM}}(x):=f\!\left(x+\rho\,u(x)\right),\quad u(x):=\frac{\nabla f(x)}{\|\nabla f(x)\|}$
+for $\nabla f(x)\neq 0$.
+However, the standard SAM implementation does not follow the surrogate gradient $\nabla f^{\mathrm{SAM}}(x)$.
+Instead, it uses the shifted gradient $\nabla f(x+\rho u(x))$, which we refer to as the SAM gradient.
+For brevity, we write
+$x^{+}:=x+\rho\,u(x).$
+Then the (full-batch) SAM iteration takes the form
+$x_{k+1}=x_{k}-\eta_{k}\,\nabla f(x_{k}^{+}),$ (1)
+so the dynamics can stall whenever $\nabla f(x^{+})=0$, regardless of whether $\nabla f(x)=0$.
+Definition 2.1 (Hallucinated minimizers).
+Fix $\rho>0$. A point $x\in\mathbb{R}^{d}$ is said to be a hallucinated minimizer at radius $\rho$ if
+$\nabla f(x^{+})=0\quad\text{and}\quad\nabla f(x)\neq 0,$
+and $x$ is a local minimizer of $f^{\mathrm{SAM}}(x)=f(x^{+})$.
+A useful relation between $\nabla f^{\mathrm{SAM}}(x)$ and the SAM gradient $\nabla f(x^{+})$.
+Assume in addition that $f$ is twice continuously differentiable on a neighborhood of $x$ and that $\nabla f(x)\neq 0$. Then $u$ is differentiable at $x$ and the chain rule gives
+$\nabla f^{\mathrm{SAM}}(x)=\bigl(I+\rho\,\nabla u(x)\bigr)^{\top}\nabla f(x^{+}).$
+If $I+\rho\nabla u(x)$ is nonsingular (equivalently, the perturbation map $x\mapsto x^{+}$ is locally invertible at $x$), then $\nabla f^{\mathrm{SAM}}(x)=0$ implies $\nabla f(x^{+})=0$.
+In particular, at such nondegenerate points, any local minimizer of $f^{\mathrm{SAM}}$ with $\nabla f(x)\neq 0$ automatically satisfies $\nabla f(x^{+})=0$.
+For brevity, when nondegeneracy is understood, we will therefore sometimes describe hallucinated minimizers simply as local minimizers of $f^{\mathrm{SAM}}$ with $\nabla f(x)\neq 0$.
+This nondegeneracy condition is used only to justify this shorthand; our existence results do not rely on it.
+Convexity rules out hallucinated minimizers.
+Hallucinated minimizers cannot arise when $f$ is convex. Indeed, Proposition A.1 in Appendix A shows that if $f$ is convex and $\nabla f(x)\neq 0$, then $\nabla f(x^{+})\neq 0$.
 $C_{\varepsilon}$$\rho=\|x_{h}-x_{\star}\|$$\nabla f(x_{h})$$x_{\star}$$_{h}xx^{\bullet}$ $x_{\star}$: global minimizer $x^{\bullet}$: local maximizer $C_{\varepsilon}$: superlevel set near $x^{\bullet}$ $x_{h}$: farthest from $x_{\star}$ on $C_{\varepsilon}$ $\nabla f(x_{h})$ points toward $x_{\star}$
@@
 By the method of Lagrange multipliers, its gradient $\nabla f(x_{h})$ points exactly toward $x_{\star}$.
+2.1 Simplified existence proof with isolated maximizers
+We now present a geometric construction that produces a hallucinated minimizer.
+For clarity, we begin with the case where $f$ admits an isolated local maximizer, namely, a point $x^{\bullet}$ with an open neighborhood $U$ such that for all $x\in U\setminus\{x^{\bullet}\}$,
+$f(x)<f(x^{\bullet})\quad\text{and}\quad\nabla f(x)\neq 0.$
+Under this assumption, the following theorem guarantees the existence of a hallucinated minimizer in a large-$\rho$ regime.
+Theorem 2.2.
+Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be continuously differentiable.
+Assume $f$ has a global minimizer $x_{\star}$ (not necessarily unique) and an isolated local maximizer $x^{\bullet}$.
+Then, there exists a nontrivial interval of radii $\rho$ for which hallucinated minimizers exist, and this interval contains values with $\rho\geq\|x_{\star}-x^{\bullet}\|$.
 Sketch of proof.
-We provide a brief sketch of the argument, with full details deferred to Appendix A.1. Figure 2 illustrates the key idea of the construction.
-[…]
-Next, consider $g(x)=\|x-x_{\star}\|^{2}$ and define
-$x_{h}\in\operatorname*{arg\,max}_{x\in C_{\varepsilon}}g(x)=\operatorname*{arg\,max}_{x\in\partial C_{\varepsilon}}g(x).$
-In words, if $x_{h}$ maximizes $g$ over $C_{\varepsilon}$, then it must lie on the boundary $\partial C_{\varepsilon}$.
-Since $\partial C_{\varepsilon}$ coincides with $\{x:f(x)=f(x^{\bullet})-\varepsilon\}$ near $x_{h}$, we apply the method of Lagrange multipliers to $x_{h}\in\operatorname*{arg\,max}_{x\in\partial C_{\varepsilon}}g(x)$ to obtain
-$2(x_{\star}-x_{h})=\nabla g(x_{h})=\lambda\nabla f(x_{h}).$
-The fact that $\lambda>0$ follows from the observation that $\nabla f(x_{h})$ points “toward” both $x_{\star}$ and $x^{\bullet}$, as illustrated in Figure 2.
-Finally, setting $\rho=\lambda\|\nabla f(x_{h})\|/2$ yields
-$x_{h}+\rho\tfrac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{\star},$
-and hence $f^{\mathrm{SAM}}(x_{h})=f(x_{\star})$. Thus, $x_{h}$ is a (global) minimizer of $f^{\mathrm{SAM}}$. ∎
-Importantly, the existence of a hallucinated minimizer also holds when $x_{\star}$ is a local minimizer, provided that $f$ has a locally Lipschitz gradient. See Appendix A.2 for details.
-The proof of Theorem 2.1 reveals the conditions under which a hallucinated minimizer is likely to arise.
-Figure 2 illustrates the core idea: given a minimizer $x_{\star}$, $x_{h}$ satisfies $x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$.
+We provide a brief sketch of the argument, with full details deferred to Appendix A.1. Figure 2 illustrates the key idea.
+Let $\varepsilon>0$ and define $C_{\varepsilon}$ as the $(f(x^{\bullet})-\varepsilon)$-superlevel set restricted to a neighborhood of the isolated local maximizer $x^{\bullet}$.
+For sufficiently small $\varepsilon>0$, the set $C_{\varepsilon}$ is compact and satisfies: $(i)$ $f(x^{\bullet})-\varepsilon\leq f(x)\leq f(x^{\bullet})$ for all $x\in C_{\varepsilon}$; $(ii)$ $\nabla f$ does not vanish on $C_{\varepsilon}\setminus\{x^{\bullet}\}$; and $(iii)$ $f(x)=f(x^{\bullet})-\varepsilon$ for all $x\in\partial C_{\varepsilon}$.
+Next, consider $g(x)=\|x-x_{\star}\|^{2}$ and choose
+$x_{h}\in\operatorname*{arg\,max}_{x\in C_{\varepsilon}}g(x).$
+Then $x_{h}\in\partial C_{\varepsilon}$.
+Moreover, since $\nabla f(x_{h})\neq 0$ and $\partial C_{\varepsilon}$ locally coincides with the level set $\{x:\ f(x)=f(x^{\bullet})-\varepsilon\}$ near $x_{h}$, there exists a neighborhood $V$ of $x_{h}$ such that
+$\Sigma:=\{x\in V:\ f(x)=f(x^{\bullet})-\varepsilon\}$
+is an embedded $C^{1}$ hypersurface and $\Sigma=V\cap\partial C_{\varepsilon}$.
+Since $x_{h}$ maximizes $g$ over $C_{\varepsilon}$ and lies on $\partial C_{\varepsilon}$, it is also a local maximizer of $g$ over $\Sigma$.
+By the method of Lagrange multipliers, there exists $\lambda\in\mathbb{R}$ such that
+$\nabla g(x_{h})+\lambda\nabla f(x_{h})=0,$
+equivalently, $2(x_{\star}-x_{h})=\lambda\nabla f(x_{h})$. Furthermore, one can show that $\lambda>0$ (see Lemma A.4 in Appendix A.1).
+Let $\rho:=\|x_{\star}-x_{h}\|$. Taking norms gives $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$, and hence
+$x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{h}^{+}.$
+Therefore, $\nabla f(x_{h}^{+})=\nabla f(x_{\star})=0$, while $\nabla f(x_{h})\neq 0$ by construction.
+Moreover, $f^{\mathrm{SAM}}(x_{h})=f(x_{h}^{+})=f(x_{\star})=\min f$.
+Thus, $x_{h}$ is a (global) minimizer of $f^{\mathrm{SAM}}$ and is therefore a hallucinated minimizer.
+Finally, since $x^{\bullet}\in C_{\varepsilon}$ and $x_{h}$ maximizes $\|x-x_{\star}\|$ over $C_{\varepsilon}$, we have $\rho=\|x_{\star}-x_{h}\|\geq\|x_{\star}-x^{\bullet}\|$.
+As $\varepsilon$ varies while satisfying (i)(i)$(iii)$, the induced radius $\rho(\varepsilon):=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|$ depends continuously on $\varepsilon$ and is strictly increasing (Lemma A.5 in Appendix A.1).
+This yields an interval of radii $\rho$ for which hallucinated minimizers exist.
+∎
+Importantly, the existence of a hallucinated minimizer also holds when $x_{\star}$ is a local minimizer, provided that $f$ has a locally Lipschitz gradient; see Appendix A.2 for details.
+When hallucinated minimizers arise.
+The proof of Theorem 2.2 reveals conditions under which hallucinated minimizers are likely to arise.
+Figure 2 illustrates the core mechanism: the construction enforces $x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{h}^{+}$.
 This means that at $x_{h}$, the gradient points directly toward $x_{\star}$.
-However, near a minimizer, the gradient typically points outward, making it difficult to identify such $x_{h}$ in its immediate neighborhood.
-To resolve this, Theorem 2.1 requires the presence of a nearby maximizer, which allows the gradient to align in the desired direction and thus necessitates nonconvexity of the objective.
-The proof also shows that the perturbation radius $\rho$ must exceed the distance between minimizer and the maximizer, since $x_{h}$ is the farthest point from $x_{\star}$ on the superlevel set $C_{\varepsilon}$.
-Because $x_{h}\in C_{\varepsilon}$, a hallucinated minimizer is located near the maximizer.
-This implies that hallucinated minimizers are typically associated with high loss values.
-Taken together, these observations suggest that hallucinated minimizers generally arise in nonconvex objectives containing both local maximizers and minimizers, and that they tend to occur near a local maximizer within a high-loss region.
-2.2 Existence of hallucinated minimizers for neural networks
-Theorem 2.1 assumes that the local maximizer is isolated. We now relax this assumption, since in neural networks, maximizers (like minimizers) often occur as sets.
-To this end, we leverage the real-analyticity of neural networks, an assumption that holds when training on a finite dataset (empirical loss) with real-analytic activation functions.
-We note that all commonly used activation functions are real-analytic, except for ReLU.
-Technically, real-analyticity affords us the Łojasiewicz inequality, which we use to rule out certain pathological cases.
-Notably, our result does not rely on restrictive but common structural assumptions on the loss function, such as global smoothness or a quadratic form.
-Definition.
-Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function.
-A nonempty connected set $X^{\bullet}\subseteq\mathbb{R}^{d}$ is a local maximizer set of $f$ if there exists $\delta>0$ such that $X^{\bullet}$ is the maximizer set over its $\delta$-neighborhood.
-In other words,
-$X^{\bullet}=\operatorname*{arg\,max}_{y\in\mathcal{N}_{\delta}(X^{\bullet})}f(y).$
-Furthermore, we denote by $f(X^{\bullet})$ the common function value of $f$ on $X^{\bullet}$.
-Figure 3: Illustration of a local maximizer set. The dotted unit circle is a local maximizer set of a real-analytic function and is not a singleton.
+Near a minimizer, gradients typically point away from the minimizer, so such an alignment is hard to obtain locally.
+The presence of a nearby maximizer provides a region where gradients can align in the required direction, which explains why nonconvexity is essential.
+The construction is also inherently a large-$\rho$ phenomenon: it yields radii satisfying $\rho=\|x_{\star}-x_{h}\|\geq\|x_{\star}-x^{\bullet}\|$, and the hallucinated minimizer $x_{h}$ lies near the maximizer region (inside $C_{\varepsilon}$).
+2.2 Existence with local maximizer sets
+Theorem 2.2 assumes that the local maximizer is isolated.
+In many modern objectives, however, symmetries and flat directions can produce non-isolated maximizers (e.g., a plateau or a manifold of maximizers).
+We therefore extend the existence result to the case of a local maximizer set, i.e., a connected set that maximizes $f$ within a neighborhood.
+Definition 2.3 (Local maximizer set).
+Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be continuous. A nonempty connected set $X\subseteq\mathbb{R}^{d}$ is a local maximizer set of $f$ if there exists $\delta>0$ such that
+$X=\operatorname*{arg\,max}_{y\in\mathcal{N}_{\delta}(X)}f(y).$
+In this case, $f$ is constant on $X$, and we denote this common value by $f(X)$.
+The following theorem shows that hallucinated minimizers can still arise when the maximizers are non-isolated.
 Theorem 2.4.
 Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be real-analytic.
-Assume $f$ has a global minimizer (not necessarily unique) and a bounded local maximizer set.
-Then, a hallucinated minimizer exists for some $\rho>0$.
-The proof of Theorem 2.2, fully presented in Appendix A.4, is analogous to that of Theorem 2.1, except that the isolatedness assumption is replaced by the real-analyticity assumption.
-A key technical point is to rule out the possibility that critical points accumulate densely around $X^{\bullet}$.
-This is achieved using the Łojasiewicz inequality:
-Lemma 2.3 (Łojasiewicz (1963)).
-If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is real-analytic, then for every $p\in\mathbb{R}^{d}$, there exist an open neighborhood $U$ of $p$, a constant $C>0$, and an exponent $q\in(0,1)$ such that
-$|f(p)-f(x)|^{q}\leq C\|\nabla f(x)\|\quad\textit{ for all }\,x\in U.$
-Most modern neural networks are real-analytic.
-The real-analyticity assumption on $f$ is practical in the context of deep learning. In particular, consider the empirical loss function
-$f(\theta)=\sum_{i=1}^{N}\ell(h_{\theta}(X_{i}),Y_{i}),$
-where $\theta\in\mathbb{R}^{d}$ denotes the neural network parameters.
-If the dataset is finite ($N<\infty$), $\ell(\cdot,\cdot)$ is real-analytic in its first argument (as is the case for most commonly used losses, such as cross-entropy), and $h_{\theta}$ is a neural network built from real-analytic activation functions (e.g., tanh, ELU, GELU, SiLU, swish), then $f(\theta)$ is real-analytic.
-More concretely, $h_{\theta}$ may be a finite composition of linear layers, convolution, attention, residual connections, layer normalization, batch normalization, real-analytic activation functions, the softmax function, average pooling, and dropout.
-However, $h_{\theta}$ cannot incorporate ReLU, leaky ReLU, or max-pooling, since these operations are non-smooth and therefore non-analytic.
-Further discussion and details are provided in Appendix A.3.
-3 Geometric and dynamical properties of hallucinated minimizers
-In this section, we establish a finer geometric property of hallucinated minimizers as well as a dynamical property of the SAM iterates.
-Recall that we use the notation $u(x)=\nabla f(x)/\|\nabla f(x)\|$.
-3.1 The set of hallucinated minimizers can have manifold structures
-The following theorem establishes that when the set of true minimizers has an $m$-dimensional manifold structure, the set of hallucinated minimizers inherits the same geometric structure.
-In particular, this result explains why, in Figure 1, the set of SAM minimizers appears as a curve.
+Assume $f$ has a global minimizer $x_{\star}$ (not necessarily unique) and a bounded local maximizer set $X$.
+Then, there exists a nontrivial interval of radii $\rho$ for which hallucinated minimizers exist, and this interval contains values with $\rho\geq d(x_{\star},X)$.
+The proof of Theorem 2.4 is provided in Appendix A.3.
+What changes relative to the isolated case.
+When $X$ is not isolated, critical points may in principle accumulate near $X$, which could invalidate the “regular level set” argument needed at the boundary $\partial C_{\varepsilon}$.
+Real-analyticity rules out such pathological accumulation: via the Łojasiewicz inequality, we show that any critical point sufficiently close to $X$ with function value near $f(X)$ must lie in $X$ itself (Lemma A.9 in Appendix A.3).
+This ensures $\nabla f\neq 0$ on $\partial C_{\varepsilon}$, enabling the same Lagrange-multiplier construction as in Theorem 2.2.
+As in the isolated case, varying $\varepsilon$ yields a strictly increasing, continuous radius map $\rho(\varepsilon)=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|$, hence an interval of admissible radii.
+On real-analyticity of $f$.
+The real-analyticity assumption is used to invoke the Łojasiewicz inequality and exclude pathological accumulations of critical points near $X$.
+For finite-dataset objectives, this assumption holds whenever both the loss and the network mapping are built from real-analytic primitives (e.g., Tanh/sigmoid/softplus/GELU activations with standard smooth losses); see Appendix A.4 for a self-contained sufficient condition.
+In particular, it applies directly to our full-batch MNIST experiment with Tanh activations (Section 4.1).
+While this excludes nonsmooth architectures such as ReLU and max-pooling, we empirically observe hallucinated minimizers even in ReLU-based networks (Section 4).
+3 Dynamical and geometric properties of hallucinated minimizers
+In this section, we establish a dynamical property of the SAM iterates as well as a finer geometric property of hallucinated minimizers.
+3.1 Hallucinated minimizers can be attractors
+Section 2 established that hallucinated minimizers can exist for suitable radii.
+We now show that this phenomenon is not merely an algebraic possibility: an isolated connected component of hallucinated minimizers can be locally attracting for the (full-batch) SAM update.
 Theorem 3.1.
-Suppose $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfies the assumptions of Theorem 2.2.
-Assume $\mathcal{M}\subseteq\mathrm{argmin}f$, where $\mathcal{M}\subseteq\mathbb{R}^{d}$ is a nonempty smooth $m$-dimensional manifold.
-Let $x_{h}$ be a hallucinated minimizer with a corresponding $\rho>0$ as constructed in the proof of Theorem 2.2.
-If $I+\rho\nabla u(x_{h})\in\mathbb{R}^{d\times d}$ is nonsingular, then the set of hallucinated minimizers contains a smooth manifold of dimension $m$.
-We defer the proof to Appendix B.1.
-3.2 Hallucinated minimizers are attractors
-We have established the existence of hallucinated minimizers, but does SAM actually converge to them? Recall that the SAM iteration is given by
-$x_{k+1}=x_{k}-\eta_{k}\nabla f\!\left(x_{k}+\rho\,u(x_{k})\right),\qquad k=0,1,2,\dots,$
-where $\eta_{k}>0$ denotes the step size.
-The answer is yes—hallucinated minimizers can indeed be attractors of the SAM dynamics.
-Thus, the concern about hallucinated minimizers in neural network training is not merely hypothetical.
-In Section 4, we provide empirical evidence that SAM can converge to hallucinated minimizers.
-In this subsection, we theoretically establish local convergence to hallucinated minimizers.
-Theorem 3.2.
-Suppose $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is real-analytic, and let $H\subset\mathbb{R}^{d}$ be a bounded, connected set of hallucinated minimizers of $f$ for a fixed perturbation radius $\rho>0$.
+Suppose $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is real-analytic, and let $H\subset\mathbb{R}^{d}$ be a bounded, connected set of hallucinated minimizers (for $f$ at radius $\rho$).
 Assume there exists $\delta>0$ such that the $\delta$-neighborhood of $H$ contains no minimizers of $f^{\mathrm{SAM}}$ other than those already in $H$.
 Assume further that every $x_{h}\in H$ satisfies
-$1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0,\,\,\textup{ where }\,\mathrm{Sym}(\nabla u(x_{h}))=\frac{1}{2}(\nabla u(x_{h})+\nabla u(x_{h})^{\top}).$
-If the initialization $x_{0}$ is chosen sufficiently close to $H$, then there exists a sufficiently small fixed step size $\eta_{k}=\eta>0$ such that the SAM iterates converge to $H$, in the sense that $d(x_{k},H)\to 0$.
+$1+\rho\,\lambda_{\min}\!\bigl(\mathrm{Sym}(\nabla u(x_{h}))\bigr)>0.$
+If the initialization $x_{0}$ is chosen sufficiently close to $H$, then there exists a sufficiently small fixed step size $\eta_{k}=\eta>0$ such that the SAM iterates (1) satisfy $d(x_{k},H)\to 0$.
 We defer the proof to Appendix B.1.
-4 Empirical analyses in deep learning
-In this section, we empirically validate our theory by analyzing hallucinated minimizers in deep learning.
-We show that SAM trajectories can, in practice, converge to hallucinated minimizers.
-We further demonstrate that a simple switching strategy can effectively prevent this convergence, providing a practical safeguard for SAM against this failure mode.
+Interpreting the stability condition.
+Unlike gradient descent on $f^{\mathrm{SAM}}$, the update (1) does not a priori guarantee descent of the surrogate loss $f^{\mathrm{SAM}}$.
+The condition $1+\rho\,\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{h})))>0$ ensures a local acute-angle relation between the surrogate gradient $\nabla f^{\mathrm{SAM}}(x)$ and the SAM gradient $\nabla f(x^{+})$.
+Indeed, for any $x$ with $\nabla f(x)\neq 0$ for which $u$ is differentiable, we have $\nabla f^{\mathrm{SAM}}(x)=\bigl(I+\rho\,\nabla u(x)\bigr)^{\top}\nabla f(x^{+})$.
+Therefore, with $g:=\nabla f(x^{+})$,
+$\langle\nabla f^{\mathrm{SAM}}(x),\,g\rangle=g^{\top}\!\bigl(I+\rho\,\mathrm{Sym}(\nabla u(x))\bigr)g.$
+Hence, on a neighborhood where $I+\rho\,\mathrm{Sym}(\nabla u(x))\succeq\gamma I$ for some $\gamma>0$, we have $\langle\nabla f^{\mathrm{SAM}}(x),g\rangle\geq\gamma\|g\|^{2}$.
+Combining this with $L$-smoothness of $f^{\mathrm{SAM}}$ on that neighborhood yields the descent estimate
+$f^{\mathrm{SAM}}(x-\eta g)\leq f^{\mathrm{SAM}}(x)-\eta\Bigl(\gamma-\tfrac{L\eta}{2}\Bigr)\|g\|^{2},$
+so sufficiently small steps decrease $f^{\mathrm{SAM}}$ unless $g=0$.
+The proof of Theorem 3.1 formalizes this argument near $H$ and uses the assumption that $H$ is an isolated minimizer component of $f^{\mathrm{SAM}}$ to conclude $d(x_{k},H)\to 0$.
+A simple sufficient condition.
+Directly verifying the eigenvalue condition in Theorem 3.1 can be difficult in high dimensions, so we record a convenient sufficient condition.
+At any $x$ with $\nabla f(x)\neq 0$ and $f\in C^{2}$ in a neighborhood of $x$, we have
+$\nabla u(x)=\frac{1}{\|\nabla f(x)\|}\bigl(I-u(x)u(x)^{\top}\bigr)\nabla^{2}f(x),$
+and thus $\|\nabla u(x)\|\leq\|\nabla^{2}f(x)\|/\|\nabla f(x)\|$.
+Using $\lambda_{\min}(\mathrm{Sym}(A))\geq-\|A\|$, the stability condition $1+\rho\,\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{h})))>0$ is implied by
+$\rho\,\|\nabla^{2}f(x_{h})\|<\|\nabla f(x_{h})\|.$
+This ratio-type condition is meaningful at hallucinated minimizers since $\|\nabla f(x_{h})\|>0$ by definition.
+3.2 Hallucinated minimizers can form manifolds
+We next refine the geometry of hallucinated minimizers.
+In the toy example of Figure 1, they are not isolated and instead form a curve-like set.
+The following result shows that this is not accidental: whenever the original objective has a manifold of global minimizers, the set of hallucinated minimizers can inherit the same manifold dimension.
+Theorem 3.2.
+Assume $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{2}$.
+Let $\mathcal{M}\subseteq\operatorname*{arg\,min}f$ be a nonempty embedded $m$-dimensional $C^{1}$ manifold, and fix $x_{\star}\in\mathcal{M}$.
+Suppose there exists a point $x_{h}\in\mathbb{R}^{d}$ such that
+$x_{h}^{+}=x_{\star}\quad\text{and}\quad\nabla f(x_{h})\neq 0.$
+If in addition $I+\rho\nabla u(x_{h})$ is nonsingular, then there exist neighborhoods $V$ of $x_{h}$ and $U$ of $x_{\star}$ such that the set
+$\mathcal{H}:=\{x\in V:\ x^{+}\in\mathcal{M}\cap U\}$
+is an embedded $m$-dimensional $C^{1}$ manifold. Moreover, every $x\in\mathcal{H}$ is a hallucinated minimizer at radius $\rho$.
+Its proof is provided in Appendix B.2.
+Interpretation.
+The nonsingularity of $I+\rho\nabla u(x_{h})$ exactly means that the perturbation map $\Phi_{\rho}(x):=x+\rho u(x)$ is locally invertible at $x_{h}$ (inverse function theorem).
+Thus, near $x_{h}$, the preimage $\Phi_{\rho}^{-1}(\mathcal{M})$ is a smooth manifold of the same dimension as $\mathcal{M}$.
+Connection to our existence results.
+Under the assumptions of Theorem 2.4, we may choose $x_{\star}$ to lie on any manifold component $\mathcal{M}\subseteq\operatorname*{arg\,min}f$ (when such a component exists).
+The construction in Section 2 then produces an $x_{h}$ with $x_{h}^{+}=x_{\star}$, and Theorem 3.2 explains why the resulting hallucinated minimizers can appear as curve-/surface-like sets in practice (e.g., Figure 1 and Figure 3).
+4 Experiments
 (a) Original loss $f$ around $x_{h}$
@@
 Figure 3: Visualizations of $f$ and $f^{\mathrm{SAM}}$ around the hallucinated minimizer $x_{h}$.
-Plots (a) and (b) are taken on a 2-dimensional plane defined by $x_{h}$ and two random directions.
-These show that $x_{h}$ is indeed a minimizer of $f^{\mathrm{SAM}}$ but not a stationary point of $f$.
-Plot (c) depicts $f^{\mathrm{SAM}}$ on the 2-dimensional plane containing $x_{h}$, $x_{0}$, and $x_{N}$, where $x_{0}$ is a small perturbation of $x_{h}$ and $x_{N}$ is obtained after $N=1000$ SAM steps from $x_{0}$.
+Plots (a) and (b) are taken on a two-dimensional plane defined by $x_{h}$ and two random directions.
+These show that $x_{h}$ is not stationary for $f$, while it appears as a minimizer of $f^{\mathrm{SAM}}$ on the same plane.
+Plot (c) depicts $f^{\mathrm{SAM}}$ on the two-dimensional plane containing $x_{h}$, $x_{0}$, and $x_{N}$, where $x_{0}$ is a small perturbation of $x_{h}$ and $x_{N}$ is obtained after $N=1000$ SAM steps from $x_{0}$.
 The pink horizontal line segment indicates the set of hallucinated minimizers, showing that the SAM trajectory converges back to this set.
-4.1 SAM can converge to hallucinated minimizers
-We first present an example in which SAM converges to a non-minimizer and confirm that the point is a hallucinated minimizer.
-We then examine how frequently such convergence occurs across diverse experimental settings.
-Do we really converge to hallucinated minimizers?
-We begin with a simple neural network setting and provide an example where SAM converges to a hallucinated minimizer.
-Specifically, we train a two-layer neural network with Tanh activations on MNIST (LeCun et al., 1998) using full-batch updates, yielding a smooth objective.
-The experimental setup is described in Appendix C.
-With a perturbation radius $\rho=1.8$, we train for over 20 million steps and observe that the trajectory converges to a single point $x_{h}$.
-At this point, the SAM gradient nearly vanishes ($\|\nabla f(x_{h}+\rho\,u(x_{h}))\|=4.8\times 10^{-9}$), while the original gradient remains relatively large ($\|\nabla f(x_{h})\|=0.0627$).
-Thus, $x_{h}$ is a stationary point of $f^{\mathrm{SAM}}$ but not of $f$, showing that SAM can converge to a hallucinated stationary point.
-To verify that $x_{h}$ is indeed a hallucinated minimizer, we visualize the loss landscape around it, as shown in Figure 4. Following the method of Li et al. (2018), we define the plane
-$x(\alpha,\beta)=x_{h}+\alpha u+\beta v,$
-where $u,v\in\mathbb{R}^{d}$ are orthogonal vectors of equal norm.
-In Figure 4(a), $x_{h}$ is clearly not a minimizer of $f$, whereas in Figure 4(b), on the same plane, it appears as a minimizer of $f^{\mathrm{SAM}}$.
-This demonstrates that SAM can converge to a hallucinated minimizer in neural networks.
-To investigate the geometry around $x_{h}$ in more detail, we add a small random perturbation to $x_{h}$, yielding a nearby point $x_{0}$ with $\|x_{0}-x_{h}\|=0.1$.
-Starting from $x_{0}$, we run $N=1000$ additional SAM steps to reach $x_{N}$.
-Figure 4(c) shows the SAM objective on the plane spanned by $x_{h}$, $x_{0}$, $x_{N}$.
-The SAM trajectory (blue) is projected onto this plane, while the SAM minimizers (pink) are computed within a tolerance of $10^{-9}$.
-We observe that hallucinated minimizers form a connected set, consistent with prior work on connected minimizers in neural networks (Garipov et al., 2018) and with Theorem 3.1, which guarantees that this structure is preserved.
+We empirically validate the mechanisms identified in Sections 23 on neural network objectives.
+The implementation of our experiments is available through a repository at https://anonymous.4open.science/r/SAM-can-hallucinate-minimizers-4B82/.
+Diagnostics.
+To diagnose convergence to hallucinated minimizers along a trajectory $\{x_{k}\}$, we monitor two pointwise mismatch quantities $\texttt{GradRatio}(x):=\frac{\|\nabla f(x^{+})\|}{\|\nabla f(x)\|+\epsilon}$ and $\texttt{LossGap}(x):=f(x^{+})-f(x)$, with $\epsilon>0$ a small numerical stabilizer.
+At a hallucinated minimizer $x_{h}$, we expect $\|\nabla f(x_{h}^{+})\|\approx 0$ while $\|\nabla f(x_{h})\|>0$, hence $\texttt{GradRatio}(x_{h})\ll 1$.
+Moreover, since $x_{h}^{+}$ lies near a (local) minimizer of $f$, the “ascent” perturbation can paradoxically decrease the loss, yielding $\texttt{LossGap}(x_{h})\ll 0$.
+We emphasize that in neural networks these diagnostics are evidence consistent with hallucinated minimizers, rather than a formal certification.
 $\rho=1.0$
@@
 (b) GD $\rightarrow$ SAM switching
-Figure 5: Training loss and test accuracy for SAM on MNIST with a 2-layer network using full-batch gradients.
-The top row shows the results from SAM-only training, while the bottom row shows the results from the switching strategy as described in Section 4.2.
-Final average test accuracies are 96.77%, 90.85%, 67.88%, and 13.36% for SAM-only training, compared to 97.77%, 97.73%, 97.69%, and 97.62% for the switching strategy.
-These results demonstrate that the switching strategy improves test accuracy and stabilizes outcomes across perturbation radii.
-How common are hallucinated minimizers?
-Next, we investigate convergence to hallucinated minimizers across multiple experiments, showing that this phenomenon can occur frequently in deep learning.
-Figure 5 presents SAM training outcomes in the same full-batch setting as before, with 80 distinct seeds and 100,000 iterations for perturbation radii $\rho=1.0,1.3,1.6,1.9$.
-The top row shows SAM-only results, while the bottom row corresponds to the switching strategy discussed later.
-In this subsection, however, we focus on the SAM-only setting.
-The experiments reveal that convergence to hallucinated minimizers depends on the perturbation radius $\rho$.
-At $\rho=1.0$, training consistently converges to zero loss, indicating convergence to a true minimizer of the original loss.
-At $\rho=1.3$ and $1.6$, some trajectories stabilize at nonzero-loss points.
-At $\rho=1.9$, most trajectories converge to such nonzero-loss points, indicating that SAM predominantly reaches hallucinated minimizers.
-These findings provide strong evidence that hallucinated minimizers are not rare anomalies but occur consistently in deep learning when the perturbation radius $\rho$ is large.
-This observation aligns with Theorem 2.1, whose proof requires the perturbation radius $\rho$ to exceed the distance between a minimizer and a maximizer.
-At the same time, even for the same $\rho$, trajectories may converge either to a hallucinated minimizer or to a true minimizer, consistent with Theorem 3.2, which depends on the initialization’s proximity to a hallucinated minimizer.
-In the stochastic case, we train ResNet-18 (He et al., 2016) on CIFAR-100 (Krizhevsky, 2009) with mini-batch updates, following the FSAM implementation of Li et al.
+[…]
+4.1 SAM can converge to hallucinated minimizers
+We first provide a direct verification in a smooth, full-batch setting, and then consider larger networks and stochastic settings.
+Direct verification in a smooth full-batch setting.
+We train a two-layer network with Tanh activations on MNIST using full-batch SAM.
+This setting yields a smooth (indeed, real-analytic) objective, aligning with the assumptions underlying our existence and attractor results.
+Implementation details are in Appendix C.1.
+With perturbation radius $\rho=1.8$, we run training for a long horizon until the trajectory numerically stabilizes near a point $x_{h}$.
+At this point, the shifted gradient is nearly zero while the original gradient is not: $\|\nabla f(x_{h}^{+})\|=4.8\times 10^{-9}$, and $\|\nabla f(x_{h})\|=6.27\times 10^{-2}$.
+Thus, the practical SAM update becomes (nearly) stationary at $x_{h}$ even though $x_{h}$ is not a critical point of $f$.
+To further probe the local geometry, we visualize two-dimensional loss landscapes around $x_{h}$ following Li et al.
 (2018).
-We observe a similar trend: larger perturbation radii lead to unstable training.
-Further details of the experimental settings and results are provided in Appendix C.
+Specifically, we consider planes of the form $x(\alpha,\beta)=x_{h}+\alpha d_{1}+\beta d_{2}$ where $d_{1},d_{2}$ are random orthogonal directions of matched norm.
+Figure 3 shows that $x_{h}$ is not a minimizer of $f$ on this plane, while it appears as a local minimizer of the SAM loss $f^{\mathrm{SAM}}(x)=f(x^{+})$ on the same slice.
+This is consistent with $x_{h}$ behaving as a hallucinated minimizer.
+Local attraction in practice.
+To test whether the observed point is locally attracting (rather than an artifact of a single run), we perturb $x_{h}$ by a random displacement of size $0.1$ to form $x_{0}$, run $N=1000$ additional SAM steps, and examine the resulting point $x_{N}$.
+Figure 3(c) shows that the perturbed initialization returns to the same near-minimizer region of $f^{\mathrm{SAM}}$ around $x_{h}$ within this local slice, suggesting robust local attraction, consistent with Theorem 3.1.
+Moreover, the near-minimizer region is visibly non-isolated on the slice, qualitatively aligning with the manifold mechanism in Theorem 3.2 (see also Appendix C.1).
+How common is the failure mode? (full-batch MNIST)
+We next sweep $\rho\in\{1.0,1.3,1.6,1.9\}$ and run 80 independent seeds for 100,000 iterations.
+Figure 4 reports training loss and test accuracy.
+For $\rho=1.0$, trajectories consistently reach near-zero loss.
+As $\rho$ increases, a substantial fraction of runs plateau at positive training loss and exhibit severe accuracy degradation, matching the trajectory-level signature of convergence to a spurious stationary region for the practical SAM update.
+Notably, at the same $\rho$, different seeds can yield either near-zero loss or a plateau, consistent with the basin-of-attraction viewpoint in Theorem 3.1.
+Table 1: SAM with full-batch gradients on CIFAR-10/100 (ResNet-18): metrics at the final iterate $x_{T}$ across perturbation radii $\rho$.
+Here, $x_{T}^{+}=x_{T}+\rho\,u(x_{T})$; $f(x_{T}^{+})$ is the SAM loss; and $\nabla f(x_{T}^{+})$ is the SAM gradient.
+CIFAR-10 CIFAR-100
+$\rho$ $f(x_{T})$ $f(x_{T}^{+})$ $\lVert\nabla f(x_{T})\rVert$ $\lVert\nabla f(x_{T}^{+})\rVert$ $f(x_{T})$ $f(x_{T}^{+})$ $\lVert\nabla f(x_{T})\rVert$ $\lVert\nabla f(x_{T}^{+})\rVert$
+0.0 0.001 0.001 0.006 0.006 0.009 0.009 0.058 0.058
+0.5 0.081 0.001 3.190 0.015 0.050 0.009 0.944 0.100
+1.0 0.120 0.001 3.131 0.015 0.132 0.011 4.034 0.134
+1.5 0.773 0.001 11.932 0.011 0.419 0.012 8.119 0.136
+2.0 0.219 0.002 4.033 0.021 0.828 0.014 13.722 0.123
 Figure 5: Final test accuracy for SAM-only and the switching strategy on CIFAR-100 with ResNet-18 using stochastic gradients.
@@
 Both methods achieve peak accuracy at $\rho=0.3$, with 77.05% for SAM-only and 77.49% for the switching strategy.
-4.2 Switching strategy for avoiding hallucinated minimizers
-One obvious approach to avoiding hallucinated minimizers is to use a small perturbation radius $\rho>0$.
-Indeed, hallucinated minimizers do not arise when $\rho=0$ (i.e., when SAM is not applied).
-However, the perturbation itself is the key mechanism that regularizes against sharpness, making it desirable to use a moderately large value of $\rho$.
-In this subsection, we introduce a simple yet effective heuristic for avoiding hallucinated minimizers, which we call switching.
-The idea is to use plain gradient descent for the first 10% of training iterations and then switch to SAM.
-As shown in Figure 5(b), the switching strategy consistently drives the training loss to zero across all tested perturbation radii, including large values of $\rho$ for which standard SAM fails to converge to the true minima.
-Figure 6 further demonstrates the improved test accuracy achieved under the switching strategy.
-Not only does switching yield higher test accuracy, but it also reduces sensitivity to the choice of $\rho$.
-This is a notable advantage, since the perturbation radius $\rho$ is a difficult hyperparameter to tune in SAM.
-The effectiveness of switching can be understood within our theoretical framework.
-The proof of Theorem 2.1 suggests that hallucinated minimizers tend to arise “near” local maximizers, and hence are more likely to occur in high-loss regions.
-This further implies that during the early stages of training, when loss values are large, SAM is particularly at risk of entering a basin of attraction of hallucinated minimizers.
-By first applying gradient descent to escape from high-loss regions before switching to SAM, this risk is mitigated.
-Consequently, our simple remedy ensures that SAM remains stable even for large perturbation radii.
-A related mechanism was studied by Zhou et al.
-(2025), who applied SAM only in the later stages of training to improve generalization.
-In contrast, our results show that the switching strategy specifically prevents convergence to hallucinated minimizers, providing a complementary explanation for its effectiveness.
+Larger networks: full-batch ResNet-18 on CIFAR.
+To test whether the same mismatch arises beyond small smooth networks, we run full-batch SAM on ResNet-18 for CIFAR-10/100 across $\rho\in\{0.0,0.5,1.0,1.5,2.0\}$ (Appendix C.2).
+Although ReLU networks fall outside our smooth/analytic assumptions, the same shifted-vs-original mismatch signatures are visible.
+Table 1 reports final-step metrics and shows that for $\rho>0$, the SAM loss $f(x_{T}^{+})$ and the shifted gradient norm $\|\nabla f(x_{T}^{+})\|$ can be orders of magnitude smaller than their unshifted counterparts.
+The corresponding trajectories (Appendix C.2) plateau at strictly positive $f(x_{T})$ while $f(x_{T}^{+})$ remains small, consistent with the hallucinated-minimizer mechanism.
+Stochastic training on CIFAR-100.
+Finally, we examine the practical behavior under standard mini-batch training for CIFAR-100 with ResNet-18.
+Figure 5 shows that SAM-only training becomes increasingly unstable as $\rho$ grows, while a short SGD warm-start (next subsection) improves stability and reduces sensitivity to $\rho$.
+Since our theory is deterministic/full-batch, we make no formal claims in the stochastic regime; we report these results as an empirical observation that the large-$\rho$ degradation can also arise in practice.
+Additional details are provided in Appendix C.3.
+4.2 A short SGD warm-start as a practical safeguard
+A trivial way to reduce the risk of hallucinated minimizers is to keep $\rho$ very small, but this can weaken the sharpness-regularization effect that motivates SAM.
+Instead, we consider a minimal and widely compatible SGD warm-start schedule: we run the base optimizer (SGD) for a short initial phase, and then enable SAM for the remainder of training.
+We use a fixed warm-start fraction $10\%$ in all experiments and do not tune it.
+Warm-start schedules and related heuristics have been considered in prior work (e.g., Andriushchenko and Flammarion (2022)); our focus here is not to propose a new training recipe, but to show that a short warm-start is sufficient to mitigate the hallucinated-minimizer failure mode identified in Sections 23.
+Empirical effect.
+In the full-batch MNIST setting, Figure 4 shows that SAM-only training frequently plateaus at positive loss for larger radii, whereas the SGD warm-start reliably avoids these plateaus and drives training loss to (near-)zero across the same $\rho$ sweep.
+In the stochastic CIFAR-100 setting, Figure 5 similarly shows improved stability and reduced sensitivity to $\rho$.
+Connection to our theory.
+The existence proof in Theorem 2.2 constructs hallucinated minimizers via geometry near local maximizer regions, and Theorem 3.1 shows that such points can have basins of attraction under the practical SAM dynamics.
+Early in training, iterates typically lie in high-loss regions, where encountering these basins is more likely.
+A brief SGD warm-start moves the trajectory away from high-loss regions before the SAM perturbation is enabled, reducing the chance of entering a hallucinated-minimizer basin without requiring $\rho$ to be small.
 5 Conclusion
-In this work, we identify a previously unrecognized failure mode of SAM: its tendency to converge to hallucinated minimizers.
-Our theoretical analysis establishes the existence of such minimizers under practical assumptions, and our empirical results validate the theory by demonstrating that SAM can indeed converge to them.
-To address this, we present a simple switching strategy that effectively avoids hallucinated minimizers.
-Although our theoretical and empirical findings are consistent, gaps remain between the theoretical characterizations and broader empirical findings.
-These gaps open several interesting avenues for follow-up work.
-One direction is to extend our theoretical analysis to the setting where SAM employs stochastic gradients rather than full-batch gradients.
-While our experiments suggest that hallucinated minimizers also arise in the stochastic case, a more rigorous theoretical understanding is desirable.
-Another direction is to analyze, from a theoretical standpoint, how likely it is for SAM to converge to hallucinated minimizers.
-Our experiments show that convergence to such minimizers is common, whereas our current theory only guarantees convergence within a local neighborhood of these points.
-Yet another direction is to extend the analysis to other variants of SAM.
-Our current results rely on the normalization of the ascent direction, which renders the magnitude of $\nabla f(x)$ irrelevant in constructing hallucinated minimizers.
-For SAM variants that incorporate gradient magnitude in the ascent step, a modified analysis would be necessary.
-Reproducibility statement
-We have taken extensive measures to ensure reproducibility.
-Complete proofs of all theorems, together with detailed assumptions, are provided in Appendix A and Appendix B.
-Experimental setups, including datasets and hyperparameters, are described in Appendices C and D.
-The implementation of our main experiments is available through a repository at https://anonymous.4open.science/r/SAM-can-hallucinate-minimizers-4B82.
+We identified a previously under-emphasized mismatch in standard SAM: the practical update uses the shifted gradient $\nabla f(x^{+})$, rather than the surrogate gradient $\nabla f^{\mathrm{SAM}}(x)$.
+This mismatch enables a distinct large-$\rho$ failure mode in nonconvex landscapes: SAM can stall at hallucinated minimizers—local minimizers of $f^{\mathrm{SAM}}$ that are not stationary points of the original objective.
+We established geometric existence mechanisms (including extensions to local maximizer sets), showed that isolated hallucinated-minimizer components can be locally attracting under the discrete-time SAM iteration, and characterized how hallucinated minimizers can inherit manifold structure via local invertibility of the perturbation map.
+Empirically, our diagnostics on neural network objectives exhibit mismatch signatures consistent with this mechanism, and we report that a short initial SGD warm-start can serve as a simple safeguard that improves stability and reduces sensitivity to $\rho$.
+An important direction for future work is to develop analogous guarantees for stochastic mini-batch training as well as for nonsmooth architectures.
+It would also be valuable to analyze whether and how SAM variants that modify the ascent step suppress or amplify hallucinated minimizers.
 Appendix A Omitted details for Section 2
-We begin by showing that the hallucinated minimizers cannot exist when the loss function is convex.
-For ease of exposition, throughout the appendices we use the following notation: for $x\in\mathbb{R}^{d}$, let
-$x^{+}:=x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}.$
+We begin by showing that the hallucinated minimizers cannot exist when the function is convex.
 Proposition A.1.
-If the function $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex, then there is no point $x\in\mathbb{R}^{d}$ such that $\nabla f\left(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}\right)=0$.
+Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and differentiable, and fix $\rho>0$.
+Then for any $x\in\mathbb{R}^{d}$ with $\nabla f(x)\neq 0$, we have $\nabla f(x^{+})\neq 0$, where $x^{+}:=x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}$.
 Proof.
@@
 We now provide the full proof of the existence theorem.
-See 2.1
+See 2.2
 Proof.
-Let $x^{\bullet}$ be a local maximizer with $f(x^{\bullet})=M$, and let $C\subseteq U$ be a compact ball centered at $x^{\bullet}$ such that $f(x)<f(x^{\bullet})$ for all $x\in C$.
-The strict inequality follows from $x^{\bullet}$ being an isolated critical point.
+Let $x^{\bullet}$ be an isolated strict local maximizer with $f(x^{\bullet})=M$.
+Choose a closed ball $C\subseteq U$ centered at $x^{\bullet}$ such that $f(x)<f(x^{\bullet})$ for all $x\in C\setminus\{x^{\bullet}\}$, and $x^{\bullet}$ is the only critical point of $f$ in $C$.
 Define $m:=\max_{\partial C}f(x)<M$ and consider the preimage $f^{-1}([M-\varepsilon,M])$ where $0<\varepsilon<M-m$.
@@
 Then, $x_{h}$ must be on the boundary of $C_{\varepsilon}$, and thus $f(x_{h})=M-\varepsilon$.
-Furthermore, since $x^{\bullet}\neq x_{h}$ is the only critical point of $f$ in $C_{\varepsilon}$, the gradient at $x_{h}$ does not vanish.
+Furthermore, since $x^{\bullet}$ is the only critical point of $f$ in $C_{\varepsilon}$ and $x_{h}\neq x^{\bullet}$, we have $\nabla f(x_{h})\neq 0$.
 Consequently, there exists an open neighborhood $V$ of $x_{h}$ such that $\Sigma:=\{x\in V:f(x)=M-\varepsilon\}$ is an embedded $C^{1}$ hypersurface near $x_{h}$.
 By shrinking $V$ if necessary, we may assume $V\cap\partial C_{\varepsilon}=\Sigma$.
-Thus, maximizing $g(x)$ over $C_{\varepsilon}$ is locally equivalent to maximizing $g(x)$ over the hypersurface $\Sigma$.
-By the method of Lagrange multipliers, we obtain
-$\nabla g(x_{h})=\lambda\nabla f(x_{h}).$
-That is, there exists $\lambda>0$ such that
-$2(x_{\star}-x_{h})=\lambda\nabla f(x_{h})$
-by Lemma A.4. Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
+Thus, maximizing $g(x)$ over $C_{\varepsilon}$ is locally equivalent to maximizing $g(x)$ over the hypersurface $\Sigma=\{x\in V:\ f(x)=M-\varepsilon\}$ near $x_{h}$.
+Since $\nabla f(x_{h})\neq 0$, $\Sigma$ is a regular level set and hence an embedded $C^{1}$ hypersurface.
+Because $x_{h}$ maximizes $g$ over $C_{\varepsilon}$ and lies on $\partial C_{\varepsilon}$, it is also a local maximizer of $g$ over $\Sigma$.
+By the method of Lagrange multipliers, there exists $\lambda\in\mathbb{R}$ such that
+$\nabla g(x_{h})+\lambda\nabla f(x_{h})=0.$
+Moreover, $\lambda>0$ by Lemma A.4. It follows that
+$2(x_{\star}-x_{h})=\lambda\nabla f(x_{h}).$
+Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
 $x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|},$
-which implies that $x_{h}$ is a hallucinated minimizer. ∎
+which implies that $x_{h}$ is a hallucinated minimizer.
+This shows that for each $\varepsilon\in(0,M-m)$ there is a corresponding radius $\rho>0$ such that $x_{h}$ is a hallucinated minimizer.
+Since $\rho(\varepsilon)$ is strictly increasing and continuous in $\varepsilon$ by Lemma A.5, there exists an interval of radii $\rho$ as claimed.
+∎
 We now prove the three lemmas used in the proof of Theorem 2.2.
@@
 Take $x\in\partial C_{\varepsilon}$.
-If $f(x)>M-\varepsilon$, continuity of $f$ implies that there exists $r>0$ such that $f(y)>M-\varepsilon$ for all $y$ in an open ball centered at $x$ with radius $r$.
+If $f(x)>M-\varepsilon$, continuity of $f$ implies that there exists $r>0$ such that $f(y)>M-\varepsilon$ for all $y$ in an open ball $B_{r}(x)$ centered at $x$ with radius $r$.
+Since $B_{r}(x)\subset f^{-1}([M-\varepsilon,M])$ is connected and intersects $C_{\varepsilon}$, maximality of the connected component implies $B_{r}(x)\subset C_{\varepsilon}$, so $x$ would be an interior point.
 This contradicts the fact that $x$ is a boundary point of $C_{\varepsilon}$.
@@
 Then, $x_{h}$ is a maximizer of $g$ over the feasible region $\{x\in V:f(x)\geq M-\varepsilon\}$.
-Consider any feasible direction $d\in\mathbb{R}^{d}$ with $\langle\nabla f(x_{h}),d\rangle\geq 0$.
-Let $\gamma\colon(-\varepsilon,\varepsilon)\rightarrow\mathbb{R}^{d}$ be a smooth curve such that $\gamma(0)=x_{h}$, $\gamma^{\prime}(0)=d$, and $\gamma(t)\in V$.
+Take the direction $d:=\nabla f(x_{h})$, for which $\langle\nabla f(x_{h}),d\rangle=\|\nabla f(x_{h})\|^{2}>0$.
+Let $\gamma\colon(-\varepsilon,\varepsilon)\to\mathbb{R}^{d}$ be a smooth curve with $\gamma(0)=x_{h}$ and $\gamma^{\prime}(0)=d$.
 Then, $f(\gamma(t))\geq M-\varepsilon$ for sufficiently small $t>0$.
@@
 $\lim_{t\downarrow 0}\frac{d}{dt}g(\gamma(t))=\lim_{t\downarrow 0}\frac{d}{dt}\|\gamma(t)-x_{\star}\|^{2}=\lim_{t\downarrow 0}2\langle\gamma(t)-x_{\star},\gamma^{\prime}(t)\rangle=2\langle x_{h}-x_{\star},d\rangle\leq 0.$
-Taking the feasible direction $d=\nabla f(x_{h})$ yields
-$2\langle x_{h}-x_{\star},\nabla f(x_{h})\rangle=-2\lambda\|\nabla f(x_{h})\|^{2}\leq 0.$
+Since $d=\nabla f(x_{h})$,
+$2\langle x_{h}-x_{\star},\nabla f(x_{h})\rangle=-\lambda\|\nabla f(x_{h})\|^{2}\leq 0.$
 Since $\lambda\neq 0$, this inequality implies $\lambda>0$. ∎
+Lemma A.5.
+In the proof of Theorem 2.2, define
+$\rho(\varepsilon):=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|.$
+Then $\rho(\varepsilon)$ is continuous and strictly increasing for $0<\varepsilon<M-m$.
+Proof.
+Since $C_{\varepsilon}$ is compact (Lemma A.2), the maximum defining $\rho(\varepsilon)$ is attained.
+Moreover, $\|x-x_{\star}\|$ cannot attain its maximum in $\mathrm{int}(C_{\varepsilon})$.
+Therefore
+$\rho(\varepsilon)=\max_{x\in\partial C_{\varepsilon}}\|x-x_{\star}\|.$
+Strict monotonicity.
+Fix $0<\varepsilon_{1}<\varepsilon_{2}<M-m$.
+By construction of $C_{\varepsilon}$ in the proof of Theorem 2.2, $C_{\varepsilon_{1}}\subset C_{\varepsilon_{2}}$.
+Let $x_{1}\in\partial C_{\varepsilon_{1}}$ attain $\rho(\varepsilon_{1})$.
+By Lemma A.3, $f\equiv M-\varepsilon$ on $\partial C_{\varepsilon}$, so $\partial C_{\varepsilon_{1}}\cap\partial C_{\varepsilon_{2}}=\varnothing$.
+Thus $x_{1}\in\mathrm{int}(C_{\varepsilon_{2}})$, and there exists $\delta>0$ such that $B(x_{1},\delta)\subset C_{\varepsilon_{2}}$.
+Let $v:=(x_{1}-x_{\star})/\|x_{1}-x_{\star}\|$ and choose any $t\in(0,\delta)$.
+Then $x_{1}+tv\in C_{\varepsilon_{2}}$ and
+$\rho(\varepsilon_{2})\geq\|x_{1}+tv-x_{\star}\|\geq\|x_{1}-x_{\star}\|+t>\|x_{1}-x_{\star}\|=\rho(\varepsilon_{1}).$
+Thus $\rho(\varepsilon)$ is strictly increasing on (0,Mm)(0,M-m).
+Continuity.
+Fix $\varepsilon_{0}\in(0,M-m)$.
+By construction, $x^{\bullet}\in\mathrm{int}(C_{\varepsilon_{0}})$, and $x^{\bullet}$ is the only critical point of $f$ in $C$.
+Hence $\nabla f(x)\neq 0$ for all $x\in\partial C_{\varepsilon_{0}}$.
+In particular, the level value $M-\varepsilon_{0}$ is regular on $\partial C_{\varepsilon_{0}}$, so $\partial C_{\varepsilon}$ forms a $C^{1}$ hypersurface that varies continuously with $\varepsilon$ in a neighborhood of $\varepsilon_{0}$ (by the implicit function theorem, applied locally and combined via a finite cover of the compact set $\partial C_{\varepsilon_{0}}$).
+Since $C_{\varepsilon}$ is compact for all $\varepsilon\in(0,M-m)$ and
+$\rho(\varepsilon)=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|=\max_{x\in\partial C_{\varepsilon}}\|x-x_{\star}\|,$
+standard maximum-continuity arguments for continuous functions over continuously varying compact sets yield that $\rho(\varepsilon)$ is continuous at $\varepsilon_{0}$.
+As $\varepsilon_{0}$ was arbitrary, $\rho$ is continuous on (0,Mm)(0,M-m).
+∎
 A.2 Extending Theorem 2.2 to local minimizers
@@
 Corollary A.7.
-Suppose $f$ has a locally Lipschitz gradient. Then, Theorem 2.1 remains valid even when $x_{\star}$ is a local minimizer of $f$.
+Suppose $f$ has a locally Lipschitz gradient. Then Theorem 2.2 remains valid even when $x_{\star}$ is a local minimizer of $f$.
 Proof.
@@
 Now consider an open ball centered at $x_{h}$ with radius
-$r^{\star}:=\min\Big\{\frac{\delta}{1+\frac{2\rho L}{\|\nabla f(x_{h})\|}},r\Big\}.$
+$r^{\star}:=\min\left\{\frac{\delta}{1+\frac{2\rho L}{\|\nabla f(x_{h})\|}},r\right\}.$
 Then, for any $y\in B_{r^{\star}}(x_{h})$, we have
@@
 This implies $f^{\mathrm{SAM}}(y)\geq f^{\mathrm{SAM}}(x_{h})$, so $x_{h}$ is a local minimizer of $f^{\mathrm{SAM}}$, and therefore a hallucinated minimizer. ∎
-A.3 Most modern neural networks are real-analytic: Further discussion
-We now formalize the claim that the neural network $\theta\mapsto h_{\theta}(X)$ is real-analytic in its parameters $\theta$.
-By standard arguments, the following two lemmas establish that the final loss function
-$f(\theta)=\sum_{i=1}^{N}\ell(h_{\theta}(X_{i}),Y_{i})$
-is real-analytic under the assumptions on $h_{\theta}$ stated in Section 2.
-Lemma A.7.
-If $h_{\theta}\colon\mathbb{R}^{d}\to\mathbb{R}^{h}$ is a neural network with real-analytic activation functions, then $h_{\theta}$ is real-analytic as a function of $\theta$.
-Proof.
-Fix input data $x$ and let $z$ denote the hidden states, which depend on $x$ and $\theta$.
-Then, $z$, hence the entire network $h_{\theta}$, is a finite composition of the following real-analytic mappings:
-  • (Layer normalization) $z\mapsto\gamma\odot\frac{z-\mu_{L}(z)}{\sqrt{\sigma_{L}(z)+\varepsilon}}+\beta$, where $\mu_{L}{(z)}$ and $\sigma_{L}(z)$ are the per-sample mean and variance;
-  • (Batch normalization) $z\mapsto\gamma\odot\frac{z-\mu_{B}{(z)}}{\sqrt{\sigma_{B}(z)+\varepsilon}}+\beta$, where $\mu_{B}{(z)}$ and $\sigma_{B}(z)$ are the per-channel mean and variance;
-  • (Activation) $z\mapsto\sigma(z)$, where $\sigma$ is a real-analytic activation function;
-  • (Softmax) $z\mapsto\mu(z)$, where $\mu$ is the softmax function;
-  • (Average pooling) $z\mapsto Az$, where $A$ is a linear averaging operator;
-  • (Residual connection) $z\mapsto z+F_{\theta}(z)$, where $F$ is real-analytic in $\theta$;
-  • (Convolution layer) $(z,W_{i},b_{i})\mapsto W_{i}*z+b_{i}$, where * denotes the discrete convolution operator;
-  • (Dropout) $z\mapsto m\odot z$, where $m$ is a masking operator;
-  • (Linear layer) $(z,A_{i},b_{i})\mapsto A_{i}z+b_{i}$; and
-  • (Attention layer) $(z,W_{Q},W_{K},W_{V})\mapsto\mu\big(\frac{{Q}{K}^{\top}}{\sqrt{d_{\mathrm{attn}}}}\big){V}$, where $Q=zW_{Q}$, $K=zW_{K}$, $V=zW_{V}$, and $d_{\mathrm{attn}}$ is the size of the attention matrix $Q,K$.
-Since the composition of real-analytic functions is real-analytic, it follows that $h_{\theta}$ is real-analytic. ∎
-Finally, the next lemma establishes that $f$ is real-analytic. This follows directly from the fact that the composition of real-analytic functions is real-analytic; hence the proof is omitted.
-Lemma A.8.
-Let $h_{\theta}$ be a (finite) neural network constructed with linear layers, attention, convolution, layer normalization, and real-analytic activation functions (all commonly used activation functions except ReLU are real-analytic).
-Then, for a real-analytic loss function $\ell$,
-$f(\theta):=\frac{1}{N}\sum^{N}_{i=1}\ell(h_{\theta}(X_{i}),Y_{i})$
-is real-analytic as a function of $\theta$.
-A.4 Full proof of Theorem 2.2
+A.3 Full proof of Theorem 2.4
+Lemma A.8 (Łojasiewicz (1963)).
+If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is real-analytic, then for every $p\in\mathbb{R}^{d}$, there exist an open neighborhood $U$ of $p$, a constant $c>0$, and an exponent $q\in(0,1)$ such that
+$|f(p)-f(x)|^{q}\leq c\|\nabla f(x)\|\quad\text{for all }x\in U.$
 The following lemma shows that, under the real-analyticity assumption, critical points cannot accumulate around the local maximizer set $X$. The argument relies on the Łojasiewicz inequality.
 Lemma A.9.
-Suppose $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is real-analytic and $X$ is a bounded local maximizer set of $f$ with some $\delta>0$.
+Suppose $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is real-analytic and $X$ is a bounded local maximizer set of $f$ for some $\delta>0$.
 Then, there exists $\varepsilon>0$ with the following property: if $x$ is a critical point in the $\delta$-neighborhood of $X$ with $f(x)\geq f(X)-\varepsilon$, then $x\in X$.
@@
 Assume instead that $S\neq\emptyset$.
-For each point $x\in X$, Lemma 2.3 guarantees the existence of an open neighborhood $U_{x}$, a constant $C_{x}>0$, and an exponent $q_{x}\in(0,1)$ such that
+For each point $x\in X$, Lemma A.8 guarantees the existence of an open neighborhood $U_{x}$, a constant $C_{x}>0$, and an exponent $q_{x}\in(0,1)$ such that
 $|f(x)-f(y)|^{q_{x}}=|f(X)-f(y)|^{q_{x}}\leq C_{x}\|\nabla f(y)\|,\,\quad\,y\in U_{x}.$
@@
 $|f(X)-f(y)|^{q_{x}}\leq C_{x}\|\nabla f(y)\|=0.$
-This implies $f(y)=f(X)$, which is a contradiction.
+This implies $f(y)=f(X)$, which contradicts $X=\operatorname*{arg\,max}_{z\in\mathcal{N}_{\delta}(X)}f(z)$ because $y\notin X$.
 Hence, $S\subseteq C:=\mathcal{N}_{\delta}(X)\setminus\bigcup_{x\in X}U_{x}$.
@@
 Finally, we use Lemma A.9 to complete the proof of Theorem 2.4.
-See 2.2
+See 2.4
 Proof.
@@
 $x_{h}\in\mathrm{argmax}_{C_{\varepsilon}}g(x),\quad x_{h}\neq x_{\star},\quad\rho:=\|x_{h}-x_{\star}\|.$
-Since the only critical points in $C_{\varepsilon}$ are those in $X$, we have $\nabla f(x_{h})\neq 0$.
+Then $x_{h}\in\partial C_{\varepsilon}$ (otherwise we could move slightly in the direction $x_{h}-x_{\star}$ while staying in $C_{\varepsilon}$ and increase $g$).
+Hence $f(x_{h})=f(X)-\varepsilon<f(X)$, so $x_{h}\notin X$, and Lemma A.9 implies $\nabla f(x_{h})\neq 0$.
 Thus, there exists an open neighborhood $V$ of $x_{h}$ such that $\Sigma:=\{x\in V:f(x)=f(X)-\varepsilon\}$ is an embedded smooth hypersurface near $x_{h}$.
@@
 Then, by the method of Lagrange multipliers, we obtain
-$\nabla g(x_{h})=\lambda\nabla f(x_{h}),$
-which implies that there exists $\lambda>0$ such that
+$\nabla g(x_{h})+\lambda\nabla f(x_{h})=0.$
+Equivalently,
 $2(x_{\star}-x_{h})=\lambda\nabla f(x_{h}).$
-The positivity of $\lambda$ follows from the same reasoning as in Lemma A.4. Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
+Moreover, $\lambda>0$ follows from the same argument as in Lemma A.4. Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
 $x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|},$
-which shows that $x_{h}$ is a hallucinated minimizer. ∎
+which shows that $x_{h}$ is a hallucinated minimizer. Moreover, just like in the proof of Theorem 2.2, the above argument applies for all
+$0<\varepsilon<\min\{\varepsilon_{1},\varepsilon_{2}\}.$
+Hence, by Lemma A.10 below, we conclude that the set of radii for which a hallucinated minimizer exists contains a nontrivial interval. ∎
+Lemma A.10 (Monotonicity/continuity of $\rho(\varepsilon)$ for maximizer-set construction).
+With $C_{\varepsilon}$ defined as the connected component of $f^{-1}([f(X)-\varepsilon,f(X)])$ containing $X$, the map $\rho(\varepsilon):=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|$ is continuous and strictly increasing for $\varepsilon\in(0,\min\{\varepsilon_{1},\varepsilon_{2}\})$.
+Proof.
+The proof is identical to Lemma A.5 and uses only compactness, nestedness $C_{\varepsilon_{1}}\subset C_{\varepsilon_{2}}$ for $\varepsilon_{1}<\varepsilon_{2}$, and connectedness. ∎
+A.4 When neural networks are real-analytic: sufficient conditions
+Theorem 2.4 assumes that the objective $f$ is real-analytic.
+This assumption is used only to invoke the Łojasiewicz inequality (Lemma A.8) and thereby rule out pathological accumulations of critical points near a local maximizer set.
+Here we record a simple sufficient condition showing that real-analyticity holds for a class of smooth (deterministic) neural network objectives.
+In particular, it covers the full-batch MNIST experiment with Tanh activations in Section 4.1.
+Throughout, we consider a finite dataset $\{(X_{i},Y_{i})\}_{i=1}^{N}$ and a deterministic network map $\theta\mapsto h_{\theta}(X_{i})$ (i.e., we exclude stochastic layers such as dropout from this discussion).
+We also assume normalization layers use the standard numerical constant $\varepsilon_{\mathrm{norm}}>0$ to avoid division by zero.
+Lemma A.11 (Real-analyticity of common smooth building blocks).
+Fix $\varepsilon_{\mathrm{norm}}>0$. The following mappings are real-analytic on their natural domains:
+  • affine/linear maps and convolutions;
+  • residual additions and average pooling;
+  • softmax and log-softmax;
+  • layer normalization and batch normalization in the standard form
+    $\mathrm{Norm}(z)=\gamma\odot\frac{z-\mu(z)}{\sqrt{\sigma(z)+\varepsilon_{\mathrm{norm}}}}+\beta,$
+    where $\mu(z)$ and $\sigma(z)$ are (per-sample or per-batch) mean/variance functions of $z$.
+Moreover, any elementwise real-analytic activation $\sigma(\cdot)$ (e.g., Tanh, sigmoid, softplus, GELU, swish) is real-analytic.
+Proof.
+Affine maps, convolutions, additions, and average pooling are polynomial/linear maps, hence real-analytic.
+Softmax and log-softmax are compositions of $\exp$, finite sums, divisions by strictly positive quantities, and $\log$ applied to strictly positive quantities, hence real-analytic.
+For normalization, note that $\mu(z)$ and $\sigma(z)$ are polynomials in the entries of $z$.
+Since $\sigma(z)\geq 0$ and $\varepsilon_{\mathrm{norm}}>0$, the quantity $\sigma(z)+\varepsilon_{\mathrm{norm}}$ is strictly positive, and the map $t\mapsto 1/\sqrt{t}$ is real-analytic on $(0,\infty)$.
+Thus $\mathrm{Norm}(z)$ is a composition of real-analytic maps, hence real-analytic.
+Finally, elementwise real-analytic activations preserve real-analyticity under composition.
+∎
+Lemma A.12.
+Assume that for each fixed input $X$, the network output $\theta\mapsto h_{\theta}(X)$ is obtained by composing the building blocks in Lemma A.11 with real-analytic activations.
+Let the loss $\ell(\cdot,Y)$ be real-analytic in its first argument for each label $Y$ (e.g., squared loss; cross-entropy combined with softmax).
+Then the finite-dataset objective
+$f(\theta):=\frac{1}{N}\sum^{N}_{i=1}\ell(h_{\theta}(X_{i}),Y_{i})$
+is real-analytic in $\theta$.
+Proof.
+By Lemma A.11 and closure of real-analytic functions under composition, each map $\theta\mapsto h_{\theta}(X_{i})$ is real-analytic.
+Composing with the real-analytic loss $\ell(\cdot,Y_{i})$ preserves real-analyticity, and finite sums preserve real-analyticity.
+Hence $f$ is real-analytic.
+∎
 Appendix B Omitted details for Section 3
-B.1 Proof of Theorem 3.1
-In this subsection, we prove Theorem 3.1. The argument relies on the implicit function theorem, which we state below.
-Theorem B.1 (Implicit function theorem, Lee (2013)).
-Let $U\subseteq\mathbb{R}^{d}\times\mathbb{R}^{d}$ be an open set, and let $(x,y)$ denote the coordinates on $U$.
-Suppose $\Phi:U\to\mathbb{R}^{d}$ is a smooth function, $(a,b)\in U$, and $c=\Phi(a,b)$.
-If the $d\times d$ matrix
-$\left(\frac{\partial\Phi^{i}}{\partial y^{j}}(a,b)\right)$
-is invertible, then there exist neighborhoods $V_{0}\subseteq\mathbb{R}^{d}$ of $a$ and $W_{0}\subseteq\mathbb{R}^{d}$ of $b$, together with a smooth function $F:V_{0}\to W_{0}$, such that $\Phi^{-1}(c)\cap(V_{0}\times W_{0})$ is the graph of $F$.
-In other words, $\Phi(x,y)=c$ for $(x,y)\in V_{0}\times W_{0}$ if and only if $y=F(x)$.
-Now we prove Theorem 3.1.
+B.1 Proof of Theorem 3.1 and further discussion
+In this subsection, we prove Theorem 3.1 and then discuss the special case of isolated hallucinated minimizers.
 See 3.1
 Proof.
-Let $x_{h}$ be a hallucinated minimizer of $f$ constructed in the proof of Theorem 2.2. Then, it satisfies
-$x_{\star}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}=x_{h}+\rho u(x_{h})$
-for some perturbation radius $\rho>0$.
-Let $V$ be an open neighborhood of $x_{h}$ on which the gradient never vanishes.
-Such a neighborhood exists because $\|\nabla f(x_{h})\|>0$ and $\nabla f$ is continuous.
-Define $F:\mathbb{R}^{d}\times V\rightarrow\mathbb{R}^{d}$ by
-$F(x,y)=y+\rho\ u(y)-x.$
-Clearly, $F(x_{\star},x_{h})=0$, and by assumption, $\frac{dF}{dy}(x_{\star},x_{h})=I+\rho\nabla u(x_{h})$ is invertible.
-By the implicit function theorem, there exist open neighborhoods $U_{0}\subseteq\mathbb{R}^{d}$ of $x_{\star}$ and $V_{0}\subseteq V$ of $x_{h}$, together with a smooth map $G:U_{0}\rightarrow V_{0}$, such that $G(x_{\star})=x_{h}$ and
-$F(x,G(x))=0\quad\forall x\in U_{0}.$
-This also implies that $G$ is also a local diffeomorphism at $x_{\star}$, since the differential is invertible at $x_{\star}$:
-$I+\frac{dF}{dy}\bigg\rvert_{y=x_{h}}\frac{dG}{dx}\bigg\rvert_{x=x_{\star}}=0\iff\frac{dG}{dx}\bigg\rvert_{x=x_{\star}}=\left(\frac{dF}{dy}\right)\bigg\rvert_{y=x_{h}}^{-1}.$
-In particular, there exists an open $U\subseteq U_{0}$ around $x_{\star}$ such that $G|_{U}:U\rightarrow G(U)$ is a diffeomorphism.
-Since $U\cap\mathcal{M}$ is an $m$-dimensional manifold (without boundary), the image $G(U\cap\mathcal{M})$ under the diffeomorphism is also an $m$-dimensional manifold.
-Finally, note that any $y\in G(U\cap\mathcal{M})$ satisfies
-$x=y+\rho\frac{\nabla f(y)}{\|\nabla f(y)\|},$
-where $x=G^{-1}(y)\in\mathcal{M}\cap U$ and hence $x\in\operatorname*{arg\,min}{f}$. Thus, $G(U\cap\mathcal{M})$ forms an $m$-dimensional manifold of hallucinated minimizers. ∎
-B.2 Proof of Theorem 3.2 and further discussion
-In this subsection, we prove Theorem 3.2 and then discuss the special case of isolated hallucinated minimizers.
-See 3.2
-Proof.
-First, we claim that the set $H$ is closed (hence compact) by construction.
-Indeed, if $x_{h}\in\bar{H}$, then the corresponding function value must equal $f^{\mathrm{SAM}}(H)=\min f^{\mathrm{SAM}}$.
-By our assumption on $H$, this implies $x_{h}\in H$.
+First, we claim that the set $H$ is closed (hence compact).
+Indeed, if $x_{h}\in\bar{H}$, then by continuity of $f^{\mathrm{SAM}}$ we have $f^{\mathrm{SAM}}(x_{h})=f^{\mathrm{SAM}}(H)=\min f^{\mathrm{SAM}}$.
+By the isolation assumption on the minimizer component, this implies $x_{h}\in H$.
 Hence $H=\bar{H}$.
-Let $\mathcal{N}_{\delta}(H)$ denote the closed $\delta$-neighborhood of $H$ from the theorem assumption.
-Since $1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0$ and $\|\nabla f(x_{h})\|>0$ for all $x_{h}\in H$, there exists an open neighborhood $W$ of $H$ such that $1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x)))>0$ and $\|\nabla f(x)\|>0$ for any $x\in W$.
+Let $\mathcal{N}_{\delta}(H)$ denote the closed $\delta$-neighborhood of $H$ from the theorem assumption. Since
+$1+\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{h})))>0$
+and $\|\nabla f(x_{h})\|>0$ for all $x_{h}\in H$, there exists an open neighborhood $W$ of $H$ such that
+$1+\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x)))>0$
+and $\|\nabla f(x)\|>0$ for any $x\in W$.
 By shrinking $\mathcal{N}_{\delta}(H)$ if necessary, we may assume $\mathcal{N}_{\delta}(H)\subseteq W$ and $f^{\mathrm{SAM}}$ is real-analytic on $\mathcal{N}_{\delta}(H)$.
@@
 $C_{\rho}:=\{x:d(x,C_{\varepsilon})\leq\rho\}.$
-Then, $C_{\rho}$ is also compact. Set
-$M:=\max_{x\in C_{\rho}}\|\nabla f(x)\|>0,\qquad L:=\max_{x\in\mathcal{N}_{\delta}(H)}\|\nabla^{2}f(x)\|>0,$
-and
-$\gamma:=\min_{x\in\mathcal{N}_{\delta}(H)}(1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x)))>0.$
-Consider the SAM update with fixed $\rho>0$ and constant step size $\eta_{k}=\eta$ chosen such that
+Then, $C_{\rho}$ is also compact. Let $L>0$ be any constant such that $f^{\mathrm{SAM}}$ is $L$-smooth on $\mathcal{N}_{\delta}(H)$, and set $M:=\max_{x\in C_{\rho}}\|\nabla f(x)\|>0$. Let
+$\gamma:=\min_{x\in\mathcal{N}_{\delta}(H)}\Bigl(1+\rho\,\lambda_{\min}(\mathrm{Sym}(\nabla u(x)))\Bigr).$
+By continuity and compactness, $\gamma>0$. Consider the SAM update with fixed $\rho>0$ and constant step size $\eta_{k}=\eta$ chosen such that
 $0<\eta<\min\Big\{\frac{\delta}{2M},\ \frac{2\gamma}{L}\Big\},$
@@
 Since
-$\nabla f^{\mathrm{SAM}}(x)=(I+\rho\nabla u(x))\nabla f(x^{+}),$
+$\nabla f^{\mathrm{SAM}}(x)=(I+\rho\nabla u(x))^{\top}\nabla f(x^{+}),$
 we obtain
@@
 $=f^{\mathrm{SAM}}(x_{k})-\eta\rho\langle\nabla u(x_{k})\nabla f(x_{k}^{+}),\nabla f(x_{k}^{+})\rangle-\eta\|\nabla f(x_{k}^{+})\|^{2}+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}$
-$\leq f^{\mathrm{SAM}}(x_{k})-\eta\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{k})))\|\nabla f(x_{k}^{+})\|^{2}-\eta\|\nabla f(x_{k}^{+})\|^{2}+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}$
-$\leq f^{\mathrm{SAM}}(x_{k})-\eta\left(1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{k})))-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}$
+$\leq f^{\mathrm{SAM}}(x_{k})-\eta\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{k})))\|\nabla f(x_{k}^{+})\|^{2}-\eta\|\nabla f(x_{k}^{+})\|^{2}+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}$
+$\leq f^{\mathrm{SAM}}(x_{k})-\eta\left(1+\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{k})))-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}$
 $\leq f^{\mathrm{SAM}}(x_{k})-\eta\left(\gamma-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}$
@@
 where the last inequality follows from $0<\eta<\frac{2\gamma}{L}$.
-Moreover, since the descent property $f^{\mathrm{SAM}}(x_{k+1})<f^{\mathrm{SAM}}(x_{k})$ holds when $\eta$ is replace by $\eta t$ for $t\in[0,1]$, the line segment from $x_{k}$ to $x_{k+1}$ lies within the sublevel set $(f^{\mathrm{SAM}})^{-1}([f_{\star},f_{\star}+\varepsilon])$.
+Moreover, since the descent property $f^{\mathrm{SAM}}(x_{k+1})<f^{\mathrm{SAM}}(x_{k})$ holds when $\eta$ is replaced by $\eta t$ for $t\in[0,1]$, the line segment from $x_{k}$ to $x_{k+1}$ lies within the sublevel set $(f^{\mathrm{SAM}})^{-1}([f_{\star},f_{\star}+\varepsilon])$.
 Thus, because $x_{k}\in C_{\varepsilon}$, we conclude $x_{k+1}\in C_{\varepsilon}$ by the connectedness of $C_{\varepsilon}$.
 This completes the induction.
-Finally, since $f^{SAM}(x_{k})$ is decreasing and bounded below, we have $\eta\left(\gamma-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|\rightarrow 0$.
+Finally, since $f^{\mathrm{SAM}}(x_{k})$ is decreasing and bounded below, we have $\eta\left(\gamma-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}\rightarrow 0$.
 Hence, $\nabla f(x_{k}^{+})\to 0$.
@@
 Suppose
-$1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0.$
+$1+\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{h})))>0.$
 Then, the SAM iterates, when initialized sufficiently close to $x_{h}$, converge to $x_{h}$ for sufficiently small fixed step size $\eta_{k}=\eta$.
@@
 Hence, the $C^{2}$ assumption on $f$ is sufficient to ensure the existence of constants $L$ and $\gamma$ as in the proof of Theorem 3.1.
-Appendix C Experimental details for SAM in deep learning
-C.1 SAM with full-batch gradients
+B.2 Proof of Theorem 3.2
+See 3.2
+Proof.
+Fix $\rho>0$ and define $\Phi_{\rho}(x):=x+\rho u(x)$ on the open set
+$\Omega:=\{x\in\mathbb{R}^{d}:\ \nabla f(x)\neq 0\}.$
+Since $f\in C^{2}$, the map $u(x)=\nabla f(x)/\|\nabla f(x)\|$ is $C^{1}$ on $\Omega$, and hence $\Phi_{\rho}$ is $C^{1}$ on $\Omega$ with Jacobian
+$D\Phi_{\rho}(x)=I+\rho\nabla u(x).$
+By assumption, $x_{h}\in\Omega$ and $D\Phi_{\rho}(x_{h})=I+\rho\nabla u(x_{h})$ is nonsingular.
+Therefore, by the inverse function theorem, there exist neighborhoods $V$ of $x_{h}$ and $U$ of $x_{\star}=\Phi_{\rho}(x_{h})$ such that
+$\Phi_{\rho}:V\to U$
+is a $C^{1}$ diffeomorphism with inverse $G:=\Phi_{\rho}^{-1}:U\to V$.
+Now set
+$\mathcal{H}:=G(\mathcal{M}\cap U)=\{x\in V:\ \Phi_{\rho}(x)\in\mathcal{M}\cap U\}.$
+Since $\mathcal{M}\cap U$ is an embedded $m$-dimensional manifold and $G$ is a diffeomorphism, $\mathcal{H}$ is an embedded $m$-dimensional manifold as well.
+(a) Original loss $f$
+(b) SAM loss $f^{\mathrm{SAM}}$
+Figure 6: Visualizations of the hallucinated minimizer $x_{h}$: (a) original loss $f$ along the line between $x_{h}$ and $x_{h}^{+}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$; (b) $f^{\mathrm{SAM}}$ over the affine plane through $x_{h}$ spanned by the directions $(x_{N}-x_{h})$ and $(x_{N}^{\prime}-x_{h})$.
+(a) Final train loss
+(b) Final test accuracy
+Figure 7: Comparison of SAM-only and the switching strategy across different perturbation radii $\rho$.
+Results are obtained using SAM with full-batch gradients over 80 seeds.
+Bold lines indicate the mean, and shaded areas represent the standard deviation.
+It remains to show that every $x\in\mathcal{H}$ is a hallucinated minimizer.
+Take any $x\in\mathcal{H}$.
+By definition, $x^{+}=\Phi_{\rho}(x)\in\mathcal{M}\subseteq\operatorname*{arg\,min}f$, so $\nabla f(x^{+})=0$ and $f(x^{+})=\min f$.
+Moreover, since $x\in V\subseteq\Omega$, we have $\nabla f(x)\neq 0$.
+Finally, $f^{\mathrm{SAM}}(x)=f(x^{+})=\min f$, and hence $x$ is a (global, hence local) minimizer of $f^{\mathrm{SAM}}$ with $\nabla f(x)\neq 0$ and $\nabla f(x^{+})=0$.
+Thus, $x$ is a hallucinated minimizer.
+∎
+Appendix C Further details on experiments
+C.1 SAM with full-batch gradients on MNIST
 In Section 4.1, we train a neural network using SAM with full-batch gradients.
@@
 Figure 6(a) presents a one-dimensional view along the line connecting $x_{h}$ and $x_{h}^{+}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$, parameterized as $x(\alpha)=(1-\alpha)x_{h}+\alpha x_{h}^{+}$.
-The plot shows that $x_{h}$ is not a minimizer of the original objective and that the surrounding loss landscape differs substantially from that around $x_{h}^{+}$.
+The plot shows that $x_{h}$ is not a minimizer of the original loss and that the surrounding loss landscape differs substantially from that around $x_{h}^{+}$.
 This demonstrates that the phenomenon of SAM converging to a hallucinated minimizer is fundamentally distinct from the case in which a saddle point becomes an attractor, which requires the surrounding quadratic structure to hold (Compagnoni et al., 2023).
-(a) Original loss $f$
-(b) SAM loss $f^{\mathrm{SAM}}$
-Figure 7: Visualizations of the hallucinated minimizer $x_{h}$: (a) original loss $f$ along the line between $x_{h}$ and $x_{h}^{+}$; (b) SAM loss $f^{\mathrm{SAM}}$ over the plane defined by $x_{h}$, $x_{N}$, and $x_{N}^{\prime}$.
 Figure 6(b) extends Figure 3(c) with an additional visualization of the SAM loss.
@@
 Applying an independent perturbation followed by the same procedure gives $x_{N}^{\prime}$.
-We then consider the plane spanned by $x_{h}$, $x_{N}$, and $x_{N}^{\prime}$, parameterized as $x(\alpha,\beta)=x_{h}+\alpha u+\beta v$ with $u=x_{N}-x_{h}$ and $v$ chosen orthogonal to $u$.
+We then consider the plane spanned by $x_{h}$, $x_{N}$, and $x_{N}^{\prime}$, parameterized as $x(\alpha,\beta)=x_{h}+\alpha d_{1}+\beta d_{2}$ with $d_{1}=x_{N}-x_{h}$ and $d_{2}$ chosen orthogonal to $d_{1}$.
 On this plane, the visualization shows that the hallucinated minimizers are not confined to a one-dimensional curve but instead extend into a two-dimensional surface-like structure.
-(a) Final train loss
-(b) Final test accuracy
-Figure 8: Comparison of SAM-only and the switching strategy across different perturbation radii $\rho$.
-Results are obtained using SAM with full-batch gradients over 80 seeds.
-Bold lines indicate the mean, and shaded areas represent the standard deviation.
 In the experiments reported in Figure 4, we investigate SAM with full-batch gradients by varying both the perturbation radius and the random seeds.
@@
 The results demonstrate that the performance of SAM is highly sensitive to the perturbation radius, whereas the switching strategy maintains stable performance even for larger values of $\rho$.
+Figure 8: Full-batch SAM on CIFAR-10 with ResNet-18: training loss $f(x)$ (left) and SAM loss $f^{\mathrm{SAM}}(x)$ (right) across perturbation radii $\rho$.
+Figure 9: Full-batch SAM on CIFAR-100 with ResNet-18: training loss $f(x)$ (left) and SAM loss $f^{\mathrm{SAM}}(x)$ (right) across perturbation radii $\rho$.
 $\rho=0.1$
@@
 Figure 10: Training loss and test accuracy curves for SAM with stochastic gradients on CIFAR-100 using ResNet-18.
-The top row shows SAM-only training, while the bottom row applies a switching strategy that runs stochastic gradient descent for the first 10% of epochs before switching to SAM.
+Thin curves show individual seeds, and the bold curve shows the mean across seeds.
+The top row shows SAM-only training, while the bottom row applies a switching strategy that runs stochastic gradient descent for the first 10% of steps before switching to SAM.
 Columns correspond to perturbation radii $\rho\in\{0.1,0.4,0.7,1.0\}$.
 Final average test accuracies are 74.29%, 76.36%, 69.64%, and 61.58% for SAM-only, compared to 74.48%, 77.22%, 73.35%, and 67.93% for the switching strategy.
-C.2 SAM with stochastic gradients
-We examine whether the phenomena observed with full-batch SAM also arise in the stochastic setting, as shown in Figure 6.
-ResNet-18 is trained on CIFAR-100 with standard data augmentations, including random cropping with padding, horizontal flipping, and Cutout (DeVries & Taylor, 2017).
+C.2 SAM with full-batch gradients on CIFAR-10/100
+This subsection reports additional trajectories for the full-batch CIFAR-10/100 experiments in Section 4.1.
+We use full-batch gradients to isolate the shifted-gradient mechanism studied in our deterministic theory.
+We train ResNet-18 from scratch on CIFAR-10/100 using full-batch SAM.
+To compute a full-batch gradient in practice, we accumulate gradients over mini-batches across the entire training set.
+The base optimizer is SGD with learning rate $10^{-3}$ and momentum $0.9$.
+We use standard per-channel normalization and no data augmentation to reduce confounding effects.
+We run 10,00010{,}000 SAM steps and sweep $\rho\in\{0.0,0.5,1.0,1.5,2.0\}$.
+Figures 89 plot the trajectories of the training loss $f(x_{k})$ and the SAM loss $f^{\mathrm{SAM}}(x_{k})=f(x_{k}^{+})$.
+Across larger $\rho$, $f(x_{k})$ often plateaus above its smallest observed values, while $f(x_{k}^{+})$ remains much smaller, consistent with the mismatch signatures in Section 4.1.
+C.3 SAM with stochastic gradients on CIFAR-100
+We examine how SAM behaves under standard mini-batch training on CIFAR-100 as the perturbation radius $\rho$ varies, as summarized in Figure 5.
+ResNet-18 is trained with standard data augmentations, including random cropping with padding, horizontal flipping, and Cutout (DeVries and Taylor, 2017).
 The mini-batch size is $64$, the learning rate $0.01$, momentum $0.9$, and weight decay $10^{-4}$.
-Training proceeds for $200$ epochs with cosine-annealed learning rates (Loshchilov & Hutter, 2016), following the practical implementation of FSAM (Li et al., 2024).
-The switching strategy applies plain stochastic gradient descent for the first $10\%$ of epochs before switching to stochastic SAM.
-Under the same setting, Figure 9 reports the CIFAR-100 results, comparing SAM-only with the switching strategy.
-Experiments are conducted for perturbation radii $\rho=0.1,0.4,0.7,1.0$ across $16$ random seeds.
-Each curve shows training loss and test accuracy over epochs, with bold lines denoting the mean across seeds and shaded regions indicating the standard deviation.
-The results show that, as in the full-batch case, SAM performance degrades with larger perturbation radii, whereas the switching strategy remains stable and robust across all settings.
+Training proceeds for $200$ epochs with cosine-annealed learning rate, following the implementation details in Li et al.
+(2024).
+We compare SAM-only against a short SGD warm-start before enabling SAM: we run plain SGD for the first $10\%$ of epochs and then continue with (stochastic) SAM for the remainder.
+Under this setting, Figure 10 reports representative loss/accuracy curves at $\rho\in\{0.1,0.4,0.7,1.0\}$ over $16$ random seeds.
+Each curve shows training loss and test accuracy over epochs, with bold lines denoting the mean across seeds.
+Overall, as $\rho$ increases, SAM-only training becomes less stable and more sensitive to the choice of $\rho$, whereas the SGD warm-start improves robustness across the same sweep.
+(a) $\rho=0.0$
+(b) $\rho=0.5$
+(c) $\rho=1.0$
+(d) $\rho=1.5$
+(e) $\rho=2.0$
+(f) $\rho=2.5$
+(g) $\rho=3.0$
+(h) $\rho=3.5$
+Figure 11: SAM loss $f^{\mathrm{SAM}}(x)$ under different perturbation radii $\rho$. The corresponding minimizers of $f^{\mathrm{SAM}}$ are shown in pink.
+(a) $\rho=0.0$
+(b) $\rho=0.5$
+(c) $\rho=1.0$
+(d) $\rho=1.5$
+(e) $\rho=2.0$
+(f) $\rho=2.5$
+(g) $\rho=3.0$
+(h) $\rho=3.5$
+Figure 12: SAM gradient field $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ for different perturbation radii $\rho$. The corresponding minimizers of $f^{\mathrm{SAM}}$ are shown in pink.
 Appendix D Two-dimensional synthetic function for visualization
-To visualize how the SAM perturbation radius $\rho$ affects the objective, we introduce the following two-dimensional synthetic function (originally illustrated in Figure 1):
+To visualize how the SAM perturbation radius $\rho$ affects the SAM loss $f^{\mathrm{SAM}}$, we introduce the following two-dimensional synthetic function (originally illustrated in Figure 1):
 $f(x,y)=0.8\exp\left(-\frac{x^{2}+y^{2}}{(2.5)^{2}}\right)\cdot W_{X}(x)-\exp\left(-(x+1.55\cos(y/1.5))^{2}\right)\cdot W_{Y}(y)+1,$
@@
 $\left\{(x,y)\in\mathbb{R}^{2}\;\middle|\;x=-1.55\cos\left(\tfrac{y}{1.5}\right),\;|y|\leq 0.6\right\}.$
-Figure 1 shows the original function $f$, the SAM objective $f^{\mathrm{SAM}}$, and its gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ at perturbation radius $\rho=2.8$.
-To further examine the effect of the perturbation radius, Figures 10 and 11 illustrate how the SAM objective $f^{\mathrm{SAM}}$ and its gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ evolve as $\rho$ varies over the range $0,0.5,\ldots,3.5$.
-In this setting, the SAM minimizers are defined as the regions where $f^{\mathrm{SAM}}$ attains its minimum values; in practice, they appear either as isolated points or as continuous curve-like structures.
-(a) $\rho=0.0$
-(b) $\rho=0.5$
-(c) $\rho=1.0$
-(d) $\rho=1.5$
-(e) $\rho=2.0$
-(f) $\rho=2.5$
-(g) $\rho=3.0$
-(h) $\rho=3.5$
-Figure 10: SAM objective $f^{\mathrm{SAM}}(x)$ under different perturbation radii $\rho$. The corresponding SAM minimizers are shown in pink.
-(a) $\rho=0.0$
-(b) $\rho=0.5$
-(c) $\rho=1.0$
-(d) $\rho=1.5$
-(e) $\rho=2.0$
-(f) $\rho=2.5$
-(g) $\rho=3.0$
-(h) $\rho=3.5$
-Figure 11: SAM gradient field $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ for different perturbation radii $\rho$. The corresponding SAM minimizers are shown in pink.
-An analysis of the SAM minimizers as a function of the perturbation radius $\rho$ reveals two distinct regimes.
+Figure 1 shows the original function $f$, the SAM loss $f^{\mathrm{SAM}}$, and the SAM gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ at perturbation radius $\rho=2.8$.
+To further examine the effect of the perturbation radius, Figures 11 and 12 illustrate how the SAM loss $f^{\mathrm{SAM}}$ and the SAM gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ evolve as $\rho$ varies over the range $0,0.5,\ldots,3.5$.
+In this setting, the minimizers of $f^{\mathrm{SAM}}$ are defined as the regions where $f^{\mathrm{SAM}}$ attains its minimum values; in practice, they appear either as isolated points or as continuous curve-like structures.
+An analysis of the minimizers of $f^{\mathrm{SAM}}$ as a function of the perturbation radius $\rho$ reveals two distinct regimes.
 For small $\rho$, the minimizers approach the critical points of $f$.
@@
 (2023).
-In contrast, for larger perturbation radii, the SAM minimizers form a curve on the right-hand side, starting near the maximizer and drifting outward as $\rho$ increases.
-This behavior is consistent with the proof of Theorem 2.1, which shows that hallucinated minimizers emerge for a sufficiently large perturbation radius $\rho$.
+In contrast, for larger perturbation radii, the minimizers of $f^{\mathrm{SAM}}$ form a curve on the right-hand side, starting near the maximizer and drifting outward as $\rho$ increases.
+This behavior is consistent with the proof of Theorem 2.2, which shows that hallucinated minimizers emerge for a large $\rho$.
 Theorem 3.2 further establishes that these minimizers preserve the dimensionality of the original minimizer manifold.
 Meanwhile, the SAM gradient field shows that these minimizers act as attractors within their neighborhood, and the conditions of Theorem 3.1 are indeed satisfied.
-Taken together, these observations show that our theory offers a full theoretical explanation of the empirical phenomena of hallucinated minimizers in this example.
+Taken together, this 2D example illustrates three qualitative regimes: (i) for small $\rho$, minimizers of $f^{\mathrm{SAM}}$ track minimizers/critical regions of $f$ (with the caveat that $f^{\mathrm{SAM}}$ is undefined exactly at critical points); (ii) as $\rho$ increases, the minimizers of $f^{\mathrm{SAM}}$ can drift toward high-loss regions near maximizers, consistent with the large-$\rho$ mechanism in Theorem 2.2; and (iii) the resulting minimizers of $f^{\mathrm{SAM}}$ can form non-isolated, curve-like sets, consistent with the manifold mechanism in Theorem 3.2.
+Moreover, the SAM gradient field visually points toward these minimizer sets, suggesting local attraction in the sense of Theorem 3.1 (the eigenvalue condition can be checked numerically in this low-dimensional setting).